Observer's Log

From PIRATE
Jump to: navigation, search

Provide a detailed log of the observing night, with the date as the first entry. You may also check the Latest News/Problems page for general technical issues of the PIRATE facility.

Add any new entries at the bottom of the wiki.


SXPA288 and S382 users: check the Observing Night Briefing page.





July-September 2016

 PIRATE relocation to Observatorio del Teide, Tenerife
Construction by Baader Planetarium
Software installation and first commissioning by Sybilla Technologies

Deneb 19:52, 7 March 2017 (GMT)S382 Team Student1 Andrew Hunt, Student2 David Greenway, Student3 Peter Aston, Student4 Maureen Nordon

Deneb 19:55, 7 March 2017 (GMT)Time 19:53 Duties : Logger - M Nordon, Operator - P Aston, Sensors - D Greenway, Data - A Hunt Deneb 20:03, 7 March 2017 (GMT)Main Target J114749.04+351335.2 Deneb 20:03, 7 March 2017 (GMT)

19:57 - Editing test star co-ordintes 6h 11m 33.6, + 35 21 17.3 (J2000) Deneb 20:03, 7 March 2017 (GMT)Should use non time based as time based did not work until the live run Deneb 20:03, 7 March 2017 (GMT)20:02 lost control Deneb 20:03, 7 March 2017 (GMT)

13/10/2016

SXPA288 User 21:04, 13 October 2016 (BST) 19:08 (UTC) Present: Darren Overman (Student 2, operator), Graeme McAteer (Student 3, log keeper), Ulrich Kolb (NDA), Alan Cayless joined 19:35

SXPA288 User 21:09, 13 October 2016 (BST) 19:23 (UTC) Dome Opened

SXPA288 User 21:09, 13 October 2016 (BST) 19:35 (UTC) Sample image M92 acquiring 30s G filter - some trailing due to wind (15m/s). Peaks too low. Resample with longer exposure.

SXPA288 User 21:09, 13 October 2016 (BST) 19:41 (UTC) Sample image M92 acquiring 90s G filter

SXPA288 User 21:09, 13 October 2016 (BST) 19:55 (UTC) Sample image M92 acquiring 90s B filter

SXPA288 User 21:09, 13 October 2016 (BST) 20:04 (UTC) Acquire 2 images M92 90s G filter

SXPA288 User 21:09, 13 October 2016 (BST) 20:06 (UTC) Acquire 2 images M92 90s B filter

SXPA288 User 21:18, 13 October 2016 (BST) 20:10 (UTC) Acquire sample M29 90s G filter - over-exposed. Resample with shorter exposure.

SXPA288 User 21:20, 13 October 2016 (BST) 20:19 (UTC) Acquire sample M29 30s G filter - over-exposed. Resample with shorter exposure

SXPA288 User 21:20, 13 October 2016 (BST) 20:20 (UTC) Acquire sample M29 30s B filter - over-exposed. Resample with shorter exposure

SXPA288 User 21:30, 13 October 2016 (BST) 20:23 (UTC) Acquire sample M29 10s G & B filter

SXPA288 User 21:30, 13 October 2016 (BST) 20:29 (UTC) Acquire 9 images of M29 10s G & B filters.

SXPA288 User 21:30, 13 October 2016 (BST) 20:31 (UTC) Dean Roberts takes over as NDA. Ulrich leaves.

SXPA288 User 21:37, 13 October 2016 (BST) 20:34 (UTC) Acquire sample M39 30s G & B filters - over-exposed. Resample with shorter exposure.

SXPA288 User 21:39, 13 October 2016 (BST) 20:36 (UTC) Acquire sample M39 10s G & B filters - over-exposed. Resample with shorter exposure.

SXPA288 User 21:40, 13 October 2016 (BST) 20:40 (UTC) Acquire sample M39 5s G & B filters - over-exposed. Resample with shorter exposure.

SXPA288 User 21:42, 13 October 2016 (BST) 20:42 (UTC) Acquire sample M39 2s G & B filters - over-exposed. Resample with shorter exposure.

SXPA288 User 21:44, 13 October 2016 (BST) 20:44 (UTC) Acquire sample M39 1s G & B filters - over-exposed.

We rechecked our 1, 2 and 5 seconds samples and have decided to sacrifice the 2 brightest stars and take 2 second exposures to get enough dimmer stars.

SXPA288 User 21:54, 13 October 2016 (BST) 20:52 (UTC) Acquire 9 images of M39 2s G & B filters.

SXPA288 User 22:16, 13 October 2016 (BST) 21:05 (UTC) Images look good. Time to play.

SXPA288 User 22:16, 13 October 2016 (BST) 21:09 (UTC) Acquire sample Andromeda 1s G, B, R & luminance.

SXPA288 User 22:16, 13 October 2016 (BST) 21:15 (UTC) Acquire further Andromeda 3s G, B, R & luminance.

SXPA288 User 22:39, 13 October 2016 (BST) 21:33 (UTC) Acquire sample M57 (Ring Nebula) G & B 30s.

SXPA288 User 22:39, 13 October 2016 (BST) 21:38 (UTC) Acquire sample M57 (Ring Nebula) R & Lum 30s.

SXPA288 User 22:44, 13 October 2016 (BST) 21:41 (UTC) Acquire sample M27 (Dumbbell Nebula) G & B filter 30s.

SXPA288 User 22:49, 13 October 2016 (BST) 21:45 (UTC) Acquire sample M27 (Dumbbell Nebula) R & Lum filter 30s.

SXPA288 User 22:49, 13 October 2016 (BST) 21:48 (UTC) Acquire sample M81 (Bode's galaxy) G & B filter. 30s - abandoned. Too low.

SXPA288 User 22:58, 13 October 2016 (BST) 21:57 (UTC) Acquire sample M76 (Little Dumbbell) G & B filters. 30s.

SXPA288 User 23:09, 13 October 2016 (BST) 22:00 (UTC) Acquire sample M76 (Little Dumbbell) R & Lum filters. 30s.

SXPA288 User 23:09, 13 October 2016 (BST) 22:04 (UTC) Acquire sample M76 (Little Dumbbell) B & G filters. 90s.

SXPA288 User 23:09, 13 October 2016 (BST) 22:07 (UTC) Acquire sample M76 (Little Dumbbell) R & Lum filters. 90s.

SXPA288 User 23:14, 13 October 2016 (BST) 22:12 (UTC) Acquire sample M27 (Dumbbell Nebula) G B, R & Lum filter 120s.

SXPA288 User 23:23, 13 October 2016 (BST) 22:22 (UTC) Control returned to NDA.

14/10/2016

NDA Stella Bradbury

Tutor 18:59, 14 October 2016 (BST) No SXPA288 students tonight. Watched telescope take calibration data at dusk (at 17:48 UTC) and noted that camera took less than 10 minutes to cool to -20 C.

Tutor 20:50, 14 October 2016 (BST) Cloud base lifted sufficiently for a few test observations (encouraged by Alan), although all-sky camera reporting only 25% star recognition overall; bright moonlight.

Tutor 21:14, 14 October 2016 (BST) NGC 6543: single 10, 30 and 60s exposures, 2x2 binning, luminance, R, G and B filters

Tutor 21:32, 14 October 2016 (BST) NGC 6946: 12 exposures as above

Tutor 21:44, 14 October 2016 (BST) Selected Observatory Manager -> Waiting Transition and released control.

Tutor 21:51, 14 October 2016 (BST) Automated observations Kolb_00_Gaia16aye recommenced.

17/10/2016

NDA Richard Busuttil

SXPA288 User 20:15, 17 October 2016 (BST) TEST

SXPA288 User 20:25, 17 October 2016 (BST) Present : Daniel Powell (Student 3-log keeper), Adam Milson (Student 4), Craig Anning (Student 2).

SXPA288 User 20:29, 17 October 2016 (BST) Cloud base sensor triggered - waiting for conditions to improve.

SXPA288 User 20:37, 17 October 2016 (BST) Adam Milson has taken control as operator.

SXPA288 User 20:43, 17 October 2016 (BST) Cloud base sensor triggered - waiting for conditions to improve.

SXPA288 User 20:50, 17 October 2016 (BST) High humidity probably due to mist- 64.7%.

SXPA288 User 20:52, 17 October 2016 (BST) Dome opened.

SXPA288 User 20:54, 17 October 2016 (BST) M29_Test_00_M29_00_Filter_B_01 20h23m56.0s 38°31'23.0" B 30.00 2016-10-17 19:23:03

SXPA288 User 20:56, 17 October 2016 (BST) M29_Test_00_M29_00_Filter_G_01 20h23m56.0s 38°31'23.0" G 30.00 2016-10-17 19:23:03

SXPA288 User 21:07, 17 October 2016 (BST) M15_Test_00_M15_00_Filter_B_01 21h29m58.0s 12°10'01.0" B 30.00 2016-10-17 19:34:33

SXPA288 User 21:07, 17 October 2016 (BST) M15_Test_00_M15_00_Filter_B_01 21h29m58.0s 12°10'01.0" B 30.00 2016-10-17 19:34:33

SXPA288 User 21:07, 17 October 2016 (BST) M15_Test_00_M15_00_Filter_G_00 21h29m58.0s 12°10'01.0" G 30.00 2016-10-17 19:34:33

SXPA288 User 21:07, 17 October 2016 (BST) M15_Test_00_M15_00_Filter_G_01 21h29m58.0s 12°10'01.0" G 30.00 2016-10-17 19:34:33

SXPA288 User 21:18, 17 October 2016 (BST) M15_Test_90_00_M15_00_Filter_B_00 21h29m58.0s 12°10'01.0" B 90.00 2016-10-17 19:45:36

SXPA288 User 21:18, 17 October 2016 (BST) M15_Test_90_00_M15_00_Filter_G_00 21h29m58.0s 12°10'01.0" G 90.00 2016-10-17 19:45:36

SXPA288 User 21:18, 17 October 2016 (BST) M29_Test_90_00_M29_00_Filter_B_00 20h23m56.0s 38°31'23.0" B 90.00 2016-10-17 19:45:55

SXPA288 User 21:18, 17 October 2016 (BST) M29_Test_90_00_M29_00_Filter_G_00 20h23m56.0s 38°31'23.0" G 90.00 2016-10-17 19:45:55

SXPA288 User 21:22, 17 October 2016 (BST) M29_Test_90_00_M29_00_Filter_B_00 20h23m56.0s 38°31'23.0" B 90.00 2016-10-17 19:45:55

SXPA288 User 21:22, 17 October 2016 (BST) M29_Test_90_00_M29_00_Filter_G_00 20h23m56.0s 38°31'23.0" G 90.00 2016-10-17 19:45:55

SXPA288 User 21:39, 17 October 2016 (BST) M15_Test_60_00_M15_00_Filter_B_00 21h29m58.0s 12°10'01.0" B 60.00 2016-10-17 20:05:57

SXPA288 User 21:39, 17 October 2016 (BST) M15_Test_60_00_M15_00_Filter_G_00 21h29m58.0s 12°10'01.0" G 60.00 2016-10-17 20:05:57

SXPA288 User 21:39, 17 October 2016 (BST) Cloud base sensor triggered - Dome Closed - waiting for conditions to improve.

SXPA288 User 22:17, 17 October 2016 (BST) Dome opened.

SXPA288 User 22:25, 17 October 2016 (BST) M29_Test_60_00_M29_00_Filter_G_00 20h23m56.0s 38°31'23.0" G 60.00 2016-10-17 20:52:57

SXPA288 User 22:25, 17 October 2016 (BST) M39_Test_60_00_M39_00_Filter_B_00 21h31m48.0s 48°25'59.0" B 60.00 2016-10-17 20:53:31

SXPA288 User 22:25, 17 October 2016 (BST) M39_Test_60_00_M39_00_Filter_G_00 21h31m48.0s 48°25'59.0" G 60.00 2016-10-17 20:53:31

SXPA288 User 22:38, 17 October 2016 (BST) Dan Powell has taken control as operator.

SXPA288 User 22:44, 17 October 2016 (BST) M31_Test_60_00_M31_00_Filter_G_00 0h42m44.0s 41°16'08.0" G 60.00 2016-10-17 21:13:06

SXPA288 User 22:46, 17 October 2016 (BST) M31_Test_60_00_M31_00_Filter_B_00 0h42m44.0s 41°16'08.0" B 60.00 2016-10-17 21:13:06

SXPA288 User 22:46, 17 October 2016 (BST) M31_Test_60_00_M31_00_Filter_G_00 0h42m44.0s 41°16'08.0" G 60.00 2016-10-17 21:13:25

SXPA288 User 22:46, 17 October 2016 (BST) M31_Test_60_00_M31_00_Filter_B_00 0h42m44.0s 41°16'08.0" B 60.00 2016-10-17 21:13:25

SXPA288 User 22:49, 17 October 2016 (BST) Craig Anning has taken control as operator.

SXPA288 User 23:19, 17 October 2016 (BST) M31_Test_120_00_M31_00_Filter_G_00 0h42m44.0s 41°16'08.0" G 60.00 2016-10-17 22:05:03

SXPA288 User 23:19, 17 October 2016 (BST) M31_Test_60_00_M31_00_Filter_B_00 0h42m44.0s 41°16'08.0" B 60.00 2016-10-17 22:07:14

SXPA288 User 23:19, 17 October 2016 (BST) M31_Test_60_00_M31_00_Filter_R_00 0h42m44.0s 41°16'08.0" R 60.00 2016-10-17 22:09:28

SXPA288 User 23:24, 17 October 2016 (BST) M29_Test_00_M29_00_Filter_G_00 0h42m44.0s 41°16'08.0" G 60.00 2016-10-17 22:14:22

SXPA288 User 23:25, 17 October 2016 (BST) M29_Test_120_00_M29_00_Filter_G_00 0h42m44.0s 41°16'08.0" G 60.00 2016-10-17 22:20:37

SXPA288 User 23:28, 17 October 2016 (BST) M29_Test_120_00_M29_00_Filter_B_00 0h42m44.0s 41°16'08.0" B 60.00 2016-10-17 22:26:17

SXPA288 User 23:35, 17 October 2016 (BST) M39_Test_120_00_M39_00_Filter_G_00 0h42m44.0s 41°16'08.0" G 60.00 2016-10-17 22:32:49

SXPA288 User 23:38, 17 October 2016 (BST) M39_Test_120_00_M39_00_Filter_B_00 0h42m44.0s 41°16'08.0" B 60.00 2016-10-17 22:35:00

SXPA288 User 23:35, 17 October 2016 (BST) M15_Test_120_00_M15_00_Filter_G_00 21h29m58.0s 12°10'01.0" G 120.00 2016-10-17 22:02:59

SXPA288 User 23:35, 17 October 2016 (BST) M15_Test_120_00_M15_00_Filter_B_00 21h29m58.0s 12°10'01.0" B 120.00 2016-10-17 22:02:59

SXPA288 User 23:46, 17 October 2016 (BST) M45_Test_120_00_M45_00_Filter_R_00 3h47m00.0s 24°07'01.0" R 120.00 2016-10-17 22:15:53

SXPA288 User 23:46, 17 October 2016 (BST) M45_Test_120_00_M45_00_Filter_G_00 3h47m00.0s 24°07'01.0" G 120.00 2016-10-17 22:15:53

SXPA288 User 23:46, 17 October 2016 (BST) M45_Test_120_00_M45_00_Filter_B_00 3h47m00.0s 24°07'01.0" B 120.00 2016-10-17 22:15:53

SXPA288 User 23:54, 17 October 2016 (BST) M31_Test_240_00_M31_00_Filter_G_00 0h42m44.0s 41°16'08.0" G 240.00 2016-10-17 22:22:20

SXPA288 User 23:54, 17 October 2016 (BST) M31_Test_240_00_M31_00_Filter_B_00 0h42m44.0s 41°16'08.0" B 240.00 2016-10-17 22:22:20

SXPA288 User 23:58, 17 October 2016 (BST) M45_Test_60_00_M45_00_Filter_B_00 3h47m00.0s 24°07'01.0" B 60.00 2016-10-17 22:28:03

SXPA288 User 00:11, 18 October 2016 (BST) Control returned to NDA. Good viewing conditions for the majority of the session. Later observations affected by the full moon. Observatory found to be intuitive and simple to use.

Richard 00:31, 18 October 2016 (BST) PIRATE observations of science targets have been manually triggered by Abot team.

18/10/2016

NDA Stella Bradbury

Tutor 19:09, 18 October 2016 (BST) Bias and Dark frames recorded. Observatory currently in cloud (cloud base approx. 2000 m, humidity 89%)

SXPA288 User 21:00, Kit Kennedy = Student 1, Murtaza Ali Khan = Student 2, Aman Kumar = Student 3 and Robbie Slattery = Student 4

SXPA288 User 21:02, Student 1 checking Image quality, Student 2 on logger, Student 3 take over the telescope, Student 4 Environmental check

SXPA288 User 21:03, 18 October 2016 (BST) Student 3 take over (cloud base still not clear)


SXPA288 User 21:16, 18 October 2016 (BST) Student 3 adding a schedule for M52, 30 sec set back exposure time, G filter, Binning 2x2, Repeat 2

SXPA288 User 21:47, 18 October 2016 (BST) Student 2 log out to swap with Student 3, Student 1 swap with Student 4

SXPA288 User 21:55, 18 October 2016 (BST) NDA pointed out that Average wind direction/ wind direction graph showing -99997.00 degrees.

SXPA288 User 22:09, 18 October 2016 (BST) Cloud Base seems to improve quickly, currently at 2753.30 m

SXPA288 User 22:13, 18 October 2016 (BST) All Sky Camera reading -Star detection is 2%

SXPA288 User 22:27, 18 October 2016 (BST) Student 2 adding a schedule for M52, 30 sec set back exposure time, B filter, Binning 2x2, Repeat 2

SXPA288 User 22:27, 18 October 2016 (BST) Student 2 released control of PIRATE and swapped for environmental monitoring with Student 1

SXPA288 User 22:31, 18 October 2016 (BST) Student 3 log out to swap with student 4

SXPA288 User 22:33, 18 October 2016 (BST) Cloud base is now acceptable

SXPA288 User 22:36, 18 October 2016 (BST) Internal camera light remotely switched on briefly to observe open dome OK.

SXPA288 User 22:37, 18 October 2016 (BST) Student 3 switch with student 1

SXPA288 User 22:39, 18 October 2016 (BST) M52_Test_120_00_M52_00_Filter_G_00 23h24m48.0s 61°35'35.0" G 30.00 2016-10-18

SXPA288 User 22:40, 18 October 2016 (BST) Telescope carrying out position move

SXPA288 User 22:45, 18 October 2016 (BST) Student 4 log out and student 2 log in

SXPA288 User 22:47, 18 October 2016 (BST) Student 3 using telescope

SXPA288 User 22:47, 18 October 2016 (BST) M39 Student 3 adding a schedule for M15, 30 sec set back exposure time, B & G filter both, Binning 2x2, Repeat 2

SXPA288 User 22:50, 18 October 2016 (BST) Student 3: SETC_00_M15_00_Filter_B_01 21h29m58.0s 12°10'01.0" B 30.00 2016-10-18 21:18:01

2016-10-18 22:24:01


SXPA288 User 22:51, 18 October 2016 (BST) Student 3: Observation: Finished ObservingHub 2016-10-18 21:50:41 2016-10-18 22:00:41 279d84a2-bb95-41f3-bcfb-92569811d9b5

SXPA288 User 22:56, 18 October 2016 (BST) Student 3: M52 Student 3 adding a schedule for M15, 10, 15, 20 and 35 sec set back exposure time, B & G filter both, Binning 2x2, Repeat 2

SXPA288 User 22:56, 18 October 2016 (BST) Student 3: SETA_2_00_M52_00_Filter_B_00 23h24m48.0s 61°35'35.0" B 15.00 2016-10-18 21:26:21

2016-10-18 22:37:01


SXPA288 User 22:56, 18 October 2016 (BST) Student 3: SETA_2_00_M52_00_Filter_B_00 23h24m48.0s 61°35'35.0" B 10.00 2016-10-18 21:26:21

2016-10-18 22:37:01


SXPA288 User 22:56, 18 October 2016 (BST) Student 3: SETA_2_00_M52_00_Filter_B_00 23h24m48.0s 61°35'35.0" B 35.00 2016-10-18 21:26:21

2016-10-18 22:37:01


SXPA288 User 22:56, 18 October 2016 (BST) Student 3: SETA_2_00_M52_00_Filter_G_00 23h24m48.0s 61°35'35.0" G 20.00 2016-10-18 21:26:21

2016-10-18 22:37:01


SXPA288 User 22:56, 18 October 2016 (BST) Student 3: SETA_2_00_M52_00_Filter_G_00 23h24m48.0s 61°35'35.0" G 15.00 2016-10-18 21:26:21

2016-10-18 22:37:01


SXPA288 User 22:56, 18 October 2016 (BST) Student 3: SETA_2_00_M52_00_Filter_G_00 23h24m48.0s 61°35'35.0" G 10.00 2016-10-18 21:26:21

2016-10-18 22:37:01


SXPA288 User 22:56, 18 October 2016 (BST) Student 3: SETA_2_00_M52_00_Filter_G_00 23h24m48.0s 61°35'35.0" G 35.00 2016-10-18 21:26:21

2016-10-18 22:37:01


SXPA288 User 22:56, 18 October 2016 (BST) Student 3: SETB_2_00_M39_00_Filter_B_00 21h31m48.0s 48°25'59.0" B 15.00 2016-10-18 21:30:39

2016-10-18 22:41:19


SXPA288 User 22:56, 18 October 2016 (BST) Student 3: SETB_2_00_M39_00_Filter_B_00 21h31m48.0s 48°25'59.0" B 20.00 2016-10-18 21:30:39

2016-10-18


SXPA288 User 11:03, 18 October 2016 (BST) Student 3: SETB_2_00_M39_00_Filter_B_00 21h31m48.0s 48°25'59.0" B 35.00 2016-10-18 21:30:39

2016-10-18 22:41:19


SXPA288 User 11:03, 18 October 2016 (BST) Student 3: SETB_2_00_M39_00_Filter_G_00 21h31m48.0s 48°25'59.0" G 10.00 2016-10-18 21:30:39

2016-10-18 22:41:19


SXPA288 User 11:03, 18 October 2016 (BST) Student 3: SETB_2_00_M39_00_Filter_G_00 21h31m48.0s 48°25'59.0" G 15.00 2016-10-18 21:30:39

2016-10-18 22:41:19


SXPA288 User 11:03, 18 October 2016 (BST) Student 3: SETB_2_00_M39_00_Filter_G_00 21h31m48.0s 48°25'59.0" G 20.00 2016-10-18 21:30:39

2016-10-18 22:41:19


SXPA288 User 11:03, 18 October 2016 (BST) Student 3: SETB_2_00_M39_00_Filter_G_00 21h31m48.0s 48°25'59.0" G 35.00 2016-10-18 21:30:39

2016-10-18 22:41:19


SXPA288 User 11:07, 18 October 2016 (BST) Student 4: cloud base = 3385.51m, Average wind direction = 101.62°, Average wind speed = 1.94m/s, Outside Humidity = 67.07%

SXPA288 User 11:10, 18 October 2016 (BST) Student 3: Student 3 adding a schedule for M15, 10, 15, 20 and 35 sec set back exposure time, B & G filter both, Binning 2x2, Repeat 2

SXPA288 User 11:10, 18 October 2016 (BST) Student 3: SETC_2_00_M15_00_Filter_G_00 21h29m58.0s 12°10'01.0" G 35.00 2016-10-18 21:36:30

2016-10-18 22:47:10


SXPA288 User 11:10, 18 October 2016 (BST) Student 3: SETC_2_00_M15_00_Filter_B_00 21h29m58.0s 12°10'01.0" B 10.00 2016-10-18 21:36:30

2016-10-18 22:47:10


SXPA288 User 11:10, 18 October 2016 (BST) Student 3: SETC_2_00_M15_00_Filter_B_00 21h29m58.0s 12°10'01.0" B 15.00 2016-10-18 21:36:30

2016-10-18 22:47:10


SXPA288 User 11:10, 18 October 2016 (BST) Student 3: SETC_2_00_M15_00_Filter_B_00 21h29m58.0s 12°10'01.0" B 20.00 2016-10-18 21:36:30

2016-10-18


SXPA288 User 11:10, 18 October 2016 (BST) Student 3: Observation finished

SXPA288 User 11:11, 18 October 2016 (BST) Student 3: M15, M39 and M52 decrease the exposure for 2 sec and 5 sec starts

SXPA288 User 11:12, 18 October 2016 (BST) Student 3: SETC_3_00_M15_00_Filter_G_00 0h00m00.0s 0°00'00.0" G 2.00 2016-10-18 21:42:18

2016-10-18 22:46:32


SXPA288 User 11:12, 18 October 2016 (BST) Student 3: SETC_3_00_M15_00_Filter_B_00 0h00m00.0s 0°00'00.0" B 5.00 2016-10-18 21:42:18

2016-10-18 22:46:32


SXPA288 User 11:12, 18 October 2016 (BST) Student 3: SETC_3_00_M15_00_Filter_B_00 0h00m00.0s 0°00'00.0" B 2.00 2016-10-18 21:42:18

2016-10-18 22:46:32


SXPA288 User 11:14, 18 October 2016 (BST) Student 3: SETC_3_00_M15_00_Filter_B_00 0h00m00.0s 0°00'00.0" B 2.00 2016-10-18 21:42:18

2016-10-18 22:46:32


SXPA288 User 11:14, 18 October 2016 (BST) Student 3: SETA_3_00_M52_00_Filter_B_00 23h24m48.0s 61°35'35.0" B 5.00 2016-10-18 21:43:51

2016-10-18 22:48:05


SXPA288 User 11:15, 18 October 2016 (BST) Student 3: SETA_3_00_M52_00_Filter_B_00 23h24m48.0s 61°35'35.0" B 2.00 2016-10-18 21:43:51

2016-10-18


SXPA288 User 11:12, 18 October 2016 (BST) Student 4: Cloud base sinking again

SXPA288 User 11:12, 18 October 2016 (BST) Student 1: Checking the data images

SXPA288 User 11:17, 18 October 2016 (BST) Student 3: SETB_3_00_M39_00_Filter_G_00 21h31m48.0s 48°25'59.0" G 2.00 2016-10-18 21:45:45

2016-10-18 22:49:59


SXPA288 User 11:17, 18 October 2016 (BST) Student 3: SETB_3_00_M39_00_Filter_G_00 21h31m48.0s 48°25'59.0" G 2.00 2016-10-18 21:45:45

2016-10-18 22:49:59


SXPA288 User 11:17, 18 October 2016 (BST) Student 3: SETB_3_00_M39_00_Filter_B_00 21h31m48.0s 48°25'59.0" B 5.00 2016-10-18 21:45:45

2016-10-18 22:49:59


SXPA288 User 11:17, 18 October 2016 (BST) Student 3: SETB_3_00_M39_00_Filter_B_00 21h31m48.0s 48°25'59.0" B 2.00 2016-10-18 21:45:45

2016-10-18 22:49:59


SXPA288 User 11:17, 18 October 2016 (BST) Student 3: Acquiring data is finished

SXPA288 User 11:19, 18 October 2016 (BST) Student 4: Cloud base = 3355.65m, Average wind speed = 2.01m/s, Average wind direction = 108.04°, Outside Humidity = 65.71%

SXPA288 User 11:19, 18 October 2016 (BST) Student 1: Checking Images data

SXPA288 User 11:27, 18 October 2016 (BST) Student 3: releasing control

SXPA288 User 11:27, 18 October 2016 (BST) Student 4: Taking control

SXPA288 User 11:28, 18 October 2016 (BST) : Student 3 is on weather and Student 4 on telescope

SXPA288 User 11:28, 18 October 2016 (BST) Student 3 : Cloud base = 3397.06 m, Average wind speed = 2.07 m/s, Average wind direction = 110.58°, Outside temperature = 9.50°C

SXPA288 User 11:31, 18 October 2016 (BST) Student 4 : M39, Exposure time = 60, Filter B & G

SXPA288 User 11:31, 18 October 2016 (BST) Student 4 : M39_00_M39_00_Filter_G_00 21h31m48.0s 48°25'59.0" G 60.00 2016-10-18 22:03:16

2016-10-18 23:07:16


SXPA288 User 11:31, 18 October 2016 (BST) Student 4 : M39_00_M39_00_Filter_B_00 21h31m48.0s 48°25'59.0" B 60.00 2016-10-18 22:03:16

2016-10-18 23:07:16


SXPA288 User 11:31, 18 October 2016 (BST) Student 4 : M15, Exposure time = 60, Filter B & G

SXPA288 User 11:35, 18 October 2016 (BST) Student 4 : M15_00_M15_00_Filter_G_00 21h29m58.0s 12°10'01.0" G 60.00 2016-10-18 22:04:11

2016-10-18 23:08:11


SXPA288 User 11:35, 18 October 2016 (BST) Student 4 : M15_00_M15_00_Filter_B_00 21h29m58.0s 12°10'01.0" B 60.00 2016-10-18 22:04:11

2016-10-18 23:08:11


SXPA288 User 11:43, 18 October 2016 (BST) Student 3 : Wind direction = 107.52° , Average wind speed = 2.41m/s, Cloud base = 3390.23m, Outside temperature = 9.45°C, Outside Humidity = 66.25%, star detection 66% overall. Bright moon due east in all-sky camera image.

SXPA288 User 11:44, 18 October 2016 (BST) Student 4 : Observation: Finished

Tutor 00:25, 19 October 2016 (BST) Took control as NDA, selected "Load work" from the Observatory Manager options. Had to log out of the dashboard and back in again in order to get the additional "..." menu options. Reported this to software developers together with eccentric wind direction measurement (which they confirmed coincided with zero windspeed). "Kolb_00_Gaia16aye" measurements now in progress under AbotVirtualObserver.

19/10/2016

SXPA288 User 21:12, 19 October 2016 (BST) Alan Cayless - logged in as Student 1

SXPA288 User 21:12, 19 October 2016 (BST) Recording video showing image acquisition

SXPA288 User 21:20, 19 October 2016 (BST) Weather clear - cloudbase 3489 m AllSky Camera showing 85% - Wind speed 1.0 m/s

SXPA288 User 20:40, 19 October 2016 (BST) M13 30 Seconds B and G 2 repeats [Note: IR left on - would not switch off]

SXPA288 User 20:50, 19 October 2016 (BST) M29 4 - 8 - 15 - 30 Seconds B and G 5 repeats

SXPA288 User 21:00, 19 October 2016 (BST) M39 4 - 8 - 15 - 30 Seconds B and G 5 repeats

SXPA288 User 21:17, 19 October 2016 (BST) M15 15 - 30 - 60 - 120 Seconds B and G 5 repeats

SXPA288 User 21:40, 19 October 2016 (BST) M13 30 Seconds B and G 2 repeats

SXPA288 User 21:50, 19 October 2016 (BST) M52 4 - 8 - 15 - 30 Seconds B and G 5 repeats.

SXPA288 User 21:55, 19 October 2016 (BST) Noted anomaly: Select Object window in Schedule template will not resolve 2-character object name e.g. "M2" - used NGC_7654 instead. Coordinates not picked up automatically - first 5 "M2" Exposures are of M52 again ** Correction - my mistake: NGC7654 _is M52 . . . M2 should be NGC7089**

SXPA288 User 22:09, 19 October 2016 (BST) M_2_A - RA and Dec entered by hand RA 21h33m30s Dec -0:49 15 - 30 - 60 - 120 Seconds B and G 5 repeats

SXPA288 User 22:28, 19 October 2016 (BST) Failed - tracking error

22:28, 19 October 2016 (BST) 22:20 M31 Andromeda 60 Seconds each R, G, B

SXPA288 User 22:39, 19 October 2016 (BST) Tried M_2_A program again - failed again

SXPA288 User 22:39, 19 October 2016 (BST) Cleared queue - Waiting transition - release control - sign out

20/10/2016

SXPA288 User 20:27, 20 October 2016 (BST) Present: Ken Mather (Student 2, Log Keeper), Laura McCafferty (Student 4, Operator), Mike Gibbs (Student 3, Data Assessor), Craig Farrell (Student 1,Weather Monitor), Anne-Katrin Klehe (NDA)

SXPA288 User 20:52, 20 October 2016 (BST) Dome Open, No rain, No clouds

SXPA288 User 20:52, 20 October 2016 (BST) Test image 1, M15 30s G and B filter, Peaks to low. Resample with longer exposure

SXPA288 User 20:52, 20 October 2016 (BST) Test image 2, M15 60s G and B filter, Error, IR left on by mistake

SXPA288 User 20:54, 20 October 2016 (BST) Telescope stopped - Cloud base too low

SXPA288 User 21:03, 20 October 2016 (BST) No action - cloud base dropping lower

SXPA288 User 21:09, 20 October 2016 (BST) Humidity at 90%, Dome closed

SXPA288 User 21:19, 20 October 2016 (BST) Student 4 has now released control

SXPA288 User 21:21, 20 October 2016 (BST) Student 3 has now been taken control

SXPA288 User 21:27, 20 October 2016 (BST) Acquiring image, M15 60s G and B filter

SXPA288 User 21:28, 20 October 2016 (BST) Dome open

SXPA288 User 21:36, 20 October 2016 (BST) Student 1 has now taken control of Observer Log

SXPA288 User 21:47, 20 October 2016 (BST) Image, M15V01 60s, G and B filter, image accessed and deemed suitable

SXPA288 User 21:49, 20 October 2016 (BST) M39, Acquiring Test Image 1, B and G filters, 60s

SXPA288 User 21:56, 20 October 2016 (BST) M39, Test Image 1, Filter B, Overexposed

SXPA288 User 21:56, 20 October 2016 (BST) M39, Test Image 1, Filter G, Overexposed

SXPA288 User 21:57, 20 October 2016 (BST) Student 3 has now released control

SXPA288 User 21:58, 20 October 2016 (BST) Student 2 has now taken control

SXPA288 User 22:03, 20 October 2016 (BST) M39, Acquiring Test Image 2, B and G filters, 30s

SXPA288 User 22:10, 20 October 2016 (BST) M39, Test Image 2, Filter B and G both Overexposed

SXPA288 User 22:14, 20 October 2016 (BST) M39, Test Image 3, Filter B and G, 10s, both over exposed

SXPA288 User 22:15, 20 October 2016 (BST) M39, Acquiring Test Image 4, Filter B and G, 2s

SXPA288 User 22:20, 20 October 2016 (BST) M39, Test Image 4, Filter B and G, 2s, unrecognisable data

SXPA288 User 22:27, 20 October 2016 (BST) M39, Acquiring Test Image 5, Filter B and G, 8s

SXPA288 User 22:34, 20 October 2016 (BST) M39, Test Image 3, retrospectively deemed suitable with a good count

SXPA288 User 22:38, 20 October 2016 (BST) M39_Glob, Filter B and G, 10s exposure, 10 images

SXPA288 User 22:43, 20 October 2016 (BST) M52, Acquiring Test Image 1, B and G Filter, 10s

SXPA288 User 22:52, 20 October 2016 (BST) M52, Acquiring Test Image 2, B and G filter, 20s

SXPA288 User 22:54, 20 October 2016 (BST) Cloud base dropped below 3500m, dome closed

SXPA288 User 22:58, 20 October 2016 (BST) Student 2 has now released control

SXPA288 User 23:21, 20 October 2016 (BST) Student 4 logged on


SXPA288 User 23:21, 20 October 2016 (BST)student 4 Dom closed as cloud base is climbing below 3500 metres

SXPA288 User 23:33, 20 October 2016 (BST) Acquiring images M52 20s B and G filters

SXPA288 User 23:38, 20 October 2016 (BST) Student 1 releasing control to NDA

21/10/2016

NDA Richard Busuttil SXPA288 User 20:18, 21 October 2016 (BST)

Observers: Evan Whakahau (Student 1) Stefan Blomskog (Student 2) Kimberly Murray (Student 3) Richard Farmer (Student 4)
PIRATE Operator: Student 4, Data Assessor: Student 3, Enviroment: Student 2, Log Keeper: Student 1

SXPA288 User 20:30, 21 October 2016 (BST) Logging issues resolved, tasks assigned, cloud base 3570m, average wind speed 7.0 m/s, All Sky Camera 93%

SXPA288 User 20:38, 21 October 2016 (BST) Test Image: M15, 30 seconds, Binning 2x2, G and B filters, x1 repeat

SXPA288 User 20:44, 21 October 2016 (BST) Test Image revealed M31 Andromeda - Wrong Scheduled Template ran, new template submitted for M15

SXPA288 User 20:56, 21 October 2016 (BST) After analysis repeat with 60 second exposure

SXPA288 User 21:16, 21 October 2016 (BST) Cloud base 3650m, AWS 7.6 m/s, All Sky Camera Camera 68%

SXPA288 User 21:16, 21 October 2016 (BST) Test Image seemed to be mirrored in X, seems to fit the our finder chart and able to find the reference star

SXPA288 User 21:21, 21 October 2016 (BST) Re run Test Image, 90 seconds, B and G filter

Switching Roles - PIRATE Operator: Student 3 , Enviroment: Student 1, Log Keeper: Student 4, Data Assessor: 2

SXPA288 User 21:34, 21 October 2016 (BST) windspeed average 6.8 m / s; cloudbase 3600m; humidity 73%; all sky camera 55%

SXPA288 User 21:42, 21 October 2016 (BST) Group agrees the 60 s exposures are acceptable

SXPA288 User 21:42, 21 October 2016 (BST) Taking science frames for M15:

Exposures   Binning   Filter
5           2         B
5           2         G

SXPA288 User 21:52, 21 October 2016 (BST) Windspeed average 7.0 m / s; cloudbase 3614m; humidity 77%; all sky camera 52%

SXPA288 User 22:00, 21 October 2016 (BST) Windspeed average 6.6 m / s; cloudbase 3739m; humidity 74%; all sky camera 50%

SXPA288 User 22:03, 21 October 2016 (BST) Trial image of M39:

 Exposures   Binning   Filter
1           2         B
1           2         G

SXPA288 User 22:08, 21 October 2016 (BST) Trial exposures very acceptable

SXPA288 User 22:08, 21 October 2016 (BST) Taking science frames for M39

  Exposures   Binning   Filter
5           2         B
5           2         G

SXPA288 User 22:12, 21 October 2016 (BST) Windspeed average 8.1 m / s; cloudbase 3772m; humidity 64.7%; all sky camera 52%

SXPA288 User 22:16, 21 October 2016 (BST)Science frames looking very acceptable

Switching roles: student 1 - quality assurance; student 2 - PIRATE operator; student 3 - log keeper; student 4 - environmental factors

SXPA288 User 22:25, 21 October 2016 (BST) Windspeed average 9 m / s; cloudbase 3893m; humidity 59%; all sky camera 54%

SXPA288 User 22:25, 21 October 2016 (BST) Trial image M52 - 30s exposure.

Exposures   Binning   Filter
1           2         B
1           2         G

SXPA288 User 22:41, 21 October 2016 (BST) Trial images deemed underexposed and the coordinates slightly off

SXPA288 User 22:41, 21 October 2016 (BST) Windspeed average 8.94 m / s; cloudbase 3779m; humidity 67%; all sky camera 92%

SXPA288 User 22:41, 21 October 2016 (BST) 2nd trial image of M52 at 40s exposure

Exposures   Binning   Filter
1           2         B
1           2         G

SXPA288 User 22:47, 21 October 2016 (BST) Despite the coordinates being input correctly the field of view seems to be slightly different. Image is deemed over exposed.

SXPA288 User 22:49, 21 October 2016 (BST) Windspeed average 8.9 m / s; cloudbase 3740m; humidity 71%; all sky camera 92%

SXPA288 User 22:49, 21 October 2016 (BST) 3rd trial image of M52 at 20.0s exposure:

Exposures   Binning   Filter
1           2         B
1           2         G

SXPA288 User 23:00, 21 October 2016 (BST) 3rd trial images named as 40s instead of 20s

SXPA288 User 23:00, 21 October 2016 (BST) Windspeed average 7.1 m / s; cloudbase 3723m; humidity 71%; all sky camera 91%

SXPA288 User 23:00, 21 October 2016 (BST) It was decided after some discussion that a 30s exposure was the best option.

SXPA288 User 23:00, 21 October 2016 (BST) Final Images of M52 taken at 30.0s:

  Exposures   Binning   Filter
5           2         B
5           2         G

SXPA288 User 23:14, 21 October 2016 (BST) Science frames acceptable

Switching roles: student 4 - quality assurance; student 1 - PIRATE operator; student 2 - log keeper; student 3 - environmental factors

SXPA288 User 23:24, 21 October 2016 (BST) windspeed average 8.10 m / s; cloudbase 3668m; humidity 59.35%; all sky camera 92%

SXPA288 User 23:24, 21 October 2016 (BST) Extra target, trying M33. Trial image M33 - 30s exposure, filters R, G, B, Luminance.

SXPA288 User 23:32, 21 October 2016 (BST) Images deemed under exposed. Doing 2nd trial with 60s exposure.

SXPA288 User 23:32, 21 October 2016 (BST) 3rd trial image of M33 at 120s.

SXPA288 User 23:38, 21 October 2016 (BST) Trying only Luminance filter at 180s.

SXPA288 User 23:45, 21 October 2016 (BST) Windspeed average 4.8 m / s; cloudbase 3520m; humidity 48.7%; all sky camera 94%

SXPA288 User 23:48, 21 October 2016 (BST) Doing 3 remaining filters (R, G, B) at 180s for M33.

  Exposures   Binning   Filter
1           2         B
1           2         G
1           2         R
1           2         Luminance

SXPA288 User 00:07, 22 October 2016 (BST) New extra target: Uranus, four filters, 1s

SXPA288 User 00:07, 22 October 2016 (BST) 2nd try on Uranus 3s

SXPA288 User 00:07, 22 October 2016 (BST) Operator releases controls, handing over to NDA

Richard 00:09, 22 October 2016 (BST) Starting PIRATE scientific acquisitions.

22/10/2016

NDA Stella Bradbury

Tutor 19:10, 22 October 2016 (BST) Automatic bias and dark frames successfully acquired however cloud is currently sweeping past the dome with large drops of water visible on the all-sky camera i.e. no dusk flats due to rain.

Student01 Charbel Eid

Student02 Adam Theiry Student03 Fraser Fulton Student04 Paul Kitson


Log updated by Fraser Fulton


SXPA288 User 20:16, 22 October 2016 (BST) NDA briefing, waiting on weather clearing. All 4 Observing students available


SXPA288 User 20:28, 22 October 2016 (BST) Cloud base 22.01 m Raining

SXPA288 User 20:34, 22 October 2016 (BST) Star detection 37% currently

SXPA288 User 21:08, 22 October 2016 (BST) PIRATE handed over to student02

SXPA288 User 21:11, 22 October 2016 (BST) Control handed over to NDA as problems handing over. Disconnection - internet problem?


SXPA288 User 22:02, 22 October 2016 (BST) Student02 has saved schedule. In control of PIRATE


SXPA288 User 22:02, 22 October 2016 (BST) Cloud base has risen fast, approx 50m every 4/5 seconds. Currently 3861 m

SXPA288 User 21:40, 22 October 2016 (BST) PIRATE back to supervised automatic mode. Rain stopped awaiting easing time clearing. Cloud base currently incrasing

SXPA288 User 21:27, 22 October 2016 (BST) Student02 is back in control of telescope

SXPA288 User 22:13, 22 October 2016 (BST) Dome opening

SXPA288 User 22:18, 22 October 2016 (BST) Data acquired. 2/3 B filters look good 1/3 G filters good, repeating G filters

SXPA288 User 22:21, 22 October 2016 (BST) Submitted B filters again

SXPA288 User 22:27, 22 October 2016 (BST) Student04 having difficulties viewing images on astrodive

SXPA288 User 22:34, 22 October 2016 (BST) Further 60s exposure of M15 taken

SXPA288 User 22:36, 22 October 2016 (BST) Log updated by Charbel Eid

SXPA288 User 22:40, 22 October 2016 (BST) Student 2 released control

SXPA288 User 22:42, 22 October 2016 (BST) Student 4 is in control of telescope

SXPA288 User 22:52, 22 October 2016 (BST) Data acquired M39 ,30s exposure, filters G, B

SXPA288 User 22:53, 22 October 2016 (BST) Cloud base 3913m

SXPA288 User 22:58, 22 October 2016 (BST) Data acquired M39 45s exposure, filters G, B

SXPA288 User 23:05, 22 October 2016 (BST) Data acquired M39 60s exposure , filters G, B

SXPA288 User 23:18, 22 October 2016 (BST) Data acquired M39 90s exposure , filters G, B

SXPA288 User 23:18, 22 October 2016 (BST) Currently the Allstar camera is showing 94% overall star detection

SXPA288 User 23:20, 22 October 2016 (BST) Data acquired M39 105s exposure , filters G, B

SXPA288 User 23:40, 22 October 2016 (BST) Student 3 is in control of telescope

SXPA288 User 23:49, 22 October 2016 (BST) Data acquired M52, 30s exposure, filters G, B

SXPA288 User 23:49, 22 October 2016 (BST) Data acquired M52, 60s exposure, filters G, B

SXPA288 User 23:49, 22 October 2016 (BST) Currently the Allstar camera is showing 94% overall star detection

SXPA288 User 23:49, 22 October 2016 (BST) Data acquired M52, 90s exposure, filters G, B

SXPA288 User 23:59, 22 October 2016 (BST) Helix nebula co-ordinates entered; RA 22 29 38.54, Dec -20 50 13.59, 60s exposure, filter luminance

SXPA288 User 00:04, 23 October 2016 (BST) Telescope stopped at 00:00

Tutor 00:18, 23 October 2016 (BST) Control returned to NDA. In order to get the menu "..." to appear I had to log out, open a *new* incognito browser window and log back in again. Students occasionally had a similar problem during the night. Selected Observatory Manager -> Waiting Transition. Cleared immediate queue which contained student's unsuccessful requests to observe the Helix Nebula.

Tutor 01:02, 23 October 2016 (BST) The telescope dome closed smoothly (as seen under IR), but the dashboard continued to indicate that the telescope was slewing. Contacted Rafal Pawlaszek via Skype PIRATE and Coast Training room. He says it looks as if there was an error in communicating with the mount. He is looking into it further and will restart the system if it is stable.

Tutor 07:06, 23 October 2016 (BST) Restart apparently successful as telescope is now completing the night's automated observations (WASP/HAT targets).

23/10/2016

SXPA288 User 20:26, 23 October 2016 (BST) Waine James (s4), Eddie Merryweather (s2) and Paul Munn (s3). Unable to observe, cloud at 3100m

SXPA288 User 20:42, 23 October 2016 (BST) Cloud still dropping, but we have set a program for M15 ready under profile s4. Program 231016_M15.

SXPA288 User 21:31. 23 October 2016 (BST) Cloudbase still low (2986m), no observing possible.

SXPA288 User 22:06. 23 October 2016 (BST) - Program for M39 set up under profile S2. Cloudbase remains under 3 km. The wait continues.

SXPA288 User 22:23. 23 October 2016 (BST) Program for M52 set up under Profile S3. Cloudbase is currently 2626m. Still unobservable.

SXPA288 User 22:23. 23 October 2016 (BST) Relinquished control to supervised automatic mode while waiting for conditions to improve.

SXPA288 User 22:50. 23 October 2016 (BST) Cloudbase still < 2500m, session abandoned.

24/10/2016

NDA: Stella Bradbury

Tutor 18:48, 24 October 2016 (BST) All sky camera is very wet but telescope is in sun as bias and dark frames are being taken at dusk. Rain sensor is still triggered but cloud base is lifting.

SXPA288 User 20:15. 24 October 2016 (BST) Sian Parry online, communicating on Skype. Awaiting communications from Darryl. Cloudbase 2500m.

SXPA288 User 20:38. 24 October 2016 (BST) Cloud base variable ~100m

SXPA288 User 20:55. 24 October 2016 (BST) Sian (student 1) Darryl (student 3). Cloud base very variable.

SXPA288 User 21:12. 24 October 2016 (BST) Cloud base variable. Wind speed currently 8.13m/s

SXPA288 User 21:18. 24 October 2016 (BST) Observatory offline due to cloud cover. Star detection 51%.

SXPA288 User 21:18. 24 October 2016 (BST) Observatory remains offline due to cloud cover. Star detection 25%. Average wind speed 6.6m/s

SXPA288 User 22:03, 24 October 2016 (BST) Observatory remains offline due to cloud cover. Star detection 11%. Average wind speed 9.01m/s. Car headlights reflected in external camera and water droplets on all sky camera.

SXPA288 User 22:18, 24 October 2016 (BST) Still offline. Cloud base appears to be lowering, now below 1000m. Star detection down to 8%.

SXPA288 User 22:26, 24 October 2016 (BST) Darryl Cooper (Student 3) has currently logged off from the observatory and is on log duty due to broadband issues.

SXPA288 User 22:44, 24 October 2016 (BST) Sian (Student 1) is programming schedule for M15. Star detection 26%. Average wind speed 9.20 m/s

SXPA288 User 22:56, 24 October 2016 (BST) Sian (Student 1) reported to Darryl (Student 3 - currently keeping log) that dome is set to open in 8 minutes.

Tutor 23:08, 24 October 2016 (BST) wiki log sheet page has become corrupted - student who was logged in no longer has edit button and html commands are visible. I logged in and couldn't see edit buttons either but went to edit tab, editied the html and saved. Now edit buttons have reappeared. Still some html left at the top of the page however.

SXPA288 User 23:13, 24 October 2016 (BST) Dome to open in 5 minutes after briefly re-setting back to 9.59. Cloud base 3358.27 m. Star detection 84%

SXPA288 User 23:15, 24 October 2016 (BST) Dome timer reset to 9.59. Cloud base 3182.12 m

SXPA288 User 23:23, 24 October 2016 (BST) M52 programming reduced to less repeats to compensate time lost

SXPA288 User 23:27, 24 October 2016 (BST) M39, M15 programming reduced. Sian and Darryl have discussed and agreed on aborting the session at 23:30 BST if cloud base is not sufficient to open dome. Cloud base 3042 m

SXPA288 User 23:36, 24 October 2016 (BST) Final Average wind speed 3.43 m/s. Cloud base 2884 m. Star detection 87%. Sian relinquished control of telescope at 23:33 BST. If cloud base was sufficient to begin the dome countdown, it would take 10 minutes for dome to open and then another 10 minutes for M52 filter which realistically would not allow time to complete data set. Session abandoned.

Tutor 23:51, 24 October 2016 (BST) Cleared immediate queue of cluster obs. and relinquished control to AbotVirtualObserver in Supervised automatic mode. Dome is still closed, IR is off, camera at -20 deg.

25/10/2016

SXPA288 User 20:17, 25 October 2016 (BST) Ste (Student 1), Peter (Student 3), Dave (Student 4) present. Cloud base is at 2275 m and Site Humidity 99%. Waiting for conditions to improve before starting.

SXPA288 User 20:24, 25 October 2016 (BST) Ste (Student 1) has taken control. Scheduling M15 tracking ahead of time for when conditions improve.

SXPA288 User 20:31, 25 October 2016 (BST) Ste (Student 1) has swapped with Dave (Student 4) due to the scheduling system not working in his browser.

SXPA288 User 21:01, 25 October 2016 (BST) Conditions continuing to hinder. Cloud base 1633 m, Site Humidity 98.45%. Star detection at 45% but fluctuating often.

SXPA288 User 21:26, 25 October 2016 (BST) Humidity is falling gradually (97.06%), cloud base at 1234 m.

Student3 has taken control stude3nt1 incharge of log

SXPA288 User 21:44, 25 October 2016 (BST) Cloud base 1070 m, humidity 97.25%, just started raining, allsky camera 21%

SXPA288 User 22:01, 25 October 2016 (BST) Student4 took over control

SXPA288 User 22:04, 25 October 2016 (BST) Cloud base 16 m, humidity 100.0%, still raining, allsky camera 51%

SXPA288 User 22:19, 25 October 2016 (BST) Student4 unable to save program using MS Edge browser has switched over to Chrome and everything seems to be working now.

SXPA288 User 22:35, 25 October 2016 (BST) Cloud base 1.0 m, humidity 100.0%, still raining, allsky camera 6%

SXPA288 User 23:04, 25 October 2016 (BST) Cloud base 0.0 m, humidity 100.0%, raining, allsky camera 0%

SXPA288 User 23:14, 25 October 2016 (BST) Student4 has released control and we have all agreed to call it a day due to the weather conditions

26/10/2016

NDA Stella Bradbury

Tutor 10:02, 26 October 2016 (BST) Turned IR light off.

Tutor 19:17, 26 October 2016 (BST) Stopped raining about 15 mins ago, cloud base too low and site humidity too high for dusk flats. Dark and bias runs acquired.

Observers: Ella-Grace Kirton (Student 1)  David Leeds (Student 2) Andrew Spratt (Student 3) Alasdair Robertson (Student 4)

SXPA288 User 23:18, 26 October 2016 (BST) Plan discussed. Decision to observe M15 followed by M39 followed by M52. Agreed upon by all participants.

Roles - PIRATE Operator: Student 1, Data Assessor: 2/3, Environment: 2/3, Log Keeper: Student 4

SXPA288 User 20:10, 26 October 2016 (BST) Student 4 (Alasdair) and Student 2 (David) logged onto to dashboard.

SXPA288 User 20:10, 26 October 2016 (BST) 2 Men inspecting the telescope with torches. Precautionary visit due to suspected smoke from dome. Too cloudy to observe at the moment anyway.

SXPA288 User 20:26, 26 October 2016 (BST) Student 1 (Ella) has now logged on. Waiting for Student 3 (Andrew) to log on. Ulrich Kolb says we have all clear to proceed from staff on Tenerife - no sign of fire.

SXPA288 User 20:34, 26 October 2016 (BST) Cannot observe. Cloud base - 3101.40m Site humidity - 88.45%

SXPA288 User 20:54, 26 October 2016 (BST) 98% Star detection. Cloud base - 3126.01m

Changed operator 21.08 - PIRATE Operator: Student 2, Data Assessor: 1/3, Environment: 1/3, Log Keeper: Student 4

SXPA288 User 21:17, 26 October 2016 (BST) "the current cloud base limit is 3200 m and seems to be the lowest possible and still safe. The site humidity is limited to 85%" - Piotr Sybilski. Cloud base is currently 3185.27m and site humidity is 87.66%.

SXPA288 User 21:56, 26 October 2016 (BST) Cloud base okay - 3278.07m Humidity still too high - 90.62%

SXPA288 User 22:13, 26 October 2016 (BST) Student 2 creating a schedule

SXPA288 User 22:19, 26 October 2016 (BST) Schedule for M15 created. Ready to submit when telescope ready.

SXPA288 User 22:19, 26 October 2016 (BST) Schedule for M39 created. Ready to submit when telescope ready.

SXPA288 User 22:23, 26 October 2016 (BST) Schedule for M52 created. Ready to submit when telescope ready.

SXPA288 User 22:25, 26 October 2016 (BST) Humidity still too high, currently 92%. Cloud base lowering.

SXPA288 User 22:32, 26 October 2016 (BST) Student 2 released control.

SXPA288 User 22:32, 26 October 2016 (BST) Debating who would like to have a go at switching the light on and off.

SXPA288 User 22:39, 26 October 2016 (BST) Student 3 (Andrew) unable to log in due to google restrictions. Connected through skype

SXPA288 User 23:14, 26 October 2016 (BST) Session ended early due to Humidity being to high

Tutor 23:25, 26 October 2016 (BST) Control returned to AbotVirtualObserver. Dome has remained closed all evening, camera at -20 deg. I.R. is off.

27/10/2016

DeanR 19:37, 27 October 2016 (BST) Opened ABOT to see PIRATE taking flats with humidity >90% and clouds rolling over the ground, then the cloud height soon dropped and ABOT closed the dome automatically. As I wanted to check inside the dome for anything out of the ordinary I logged on and turned IR light on.

Observers: Emma Norton (Student 1), Matthew Sprange (Student 2), Daniel Clarkson (Student 3), Zoe Campbell (Student 4)
PIRATE Operator: Matthew, Log Keeper: Daniel, Environment: Emma, Data Assessor: Zoe

SXPA288 User 20:24, 27 October 2016 (BST) Weather Update - Cloud Base: 1305.87 m, Wind Speed: 33.55 km/h, Outside humidity: 94.98 %; Outside Temperature: 3.61 ºC

SXPA288 User 20:26, 27 October 2016 (BST) Logging issues experienced by Students 2 and 4. Student 2 has currently resolved their issue.

SXPA288 User 20:26, 27 October 2016 (BST) Roles assigned for all 3 viewings.

SXPA288 User 20:38, 27 October 2016 (BST) Weather Update - Cloud base: 650.05 m, Wind Speed: 30.23 km/h, Outside humidity: 94.61 %, Outside Temperature: 3.54 ºC

SXPA288 User 20:44, 27 October 2016 (BST) Student 2 has taken control of PIRATE.

SXPA288 User 20:49, 27 October 2016 (BST) Student 4 resolved login issue.

SXPA288 User 20:51, 27 October 2016 (BST) AllSky Camera - Overall: 36%

SXPA288 User 20:53, 27 October 2016 (BST) Cloud base too low to open dome and commence viewing.

SXPA288 User 20:57, 27 October 2016 (BST) Weather Update - Cloud base: 982.42 m, Wind Speed: 31.26 km/h, Outside humidity: 95.52 %, Outside Temperature: 3.35 ºC

SXPA288 User 20:58, 27 October 2016 (BST) AllSky Camera - Overall: 45%

SXPA288 User 21:16, 27 October 2016 (BST) Weather Update - Cloud base: 1160.13 m, Wind Speed: 23.82 km/h, Outside humidity: 88.61 %, Outside Temperature: 3.51 ºC

SXPA288 User 21:17, 27 October 2016 (BST) AllSky Camera - Overall: 49%

SXPA288 User 21:33, 27 October 2016 (BST) Weather Update - Cloud base: 2168.76 m, Wind Speed: 33.14 km/h, Outside humidity: 84.61 %, Outside Temperature: 3.41 ºC

SXPA288 User 21:33, 27 October 2016 (BST) AllSky Camera - Overall: 53%

SXPA288 User 21:37, 27 October 2016 (BST) Cloud base looking promising at 3425 m. AllSky camera showing 55% overall, with 80% to the south. Site humidity down to 90.70 %.

SXPA288 User 21:46, 27 October 2016 (BST) Cloud base now at acceptable levels (currently at 3488.26 m). AllSky camera: Overall 61%, Site humidity: 87.85 %.

SXPA288 User 22:09, 27 October 2016 (BST) Site humidity increased to 89.43 %.

SXPA288 User 22:43, 27 October 2016 (BST) Site humidity increased to 96.20 %. AllSky Camera: Overall 20 %.

SXPA288 User 23:12, 27 October 2016 (BST) Site humidity increased to 91.98 %. AllSky Camera: Overall 45 %.

SXPA288 User 23:34, 27 October 2016 (BST) Site humidity increased to 94.21 %. AllSky Camera: Overall 35 %.

SXPA288 User 23:38, 27 October 2016 (BST) Student 2 released control of PIRATE.

SXPA288 User 23:39, 27 October 2016 (BST) Session ended due to high humidity.

DeanR 23:55, 27 October 2016 (BST) Site humidity back up to 99.33%, AllSky Camera: 2% and cloud base now down to 1976m. Looks like bad weather is rolling in again, turned IR light in dome off and signing off NDA for the night.


28/10/2016

NDA Richard Busuttil

Student observer1 : Flashing on external camera Cloud base 3860m Wind speed 32kmh Wind direction 214 degrees Outside humidity 27% Outside temperature 9.3 degrees centugrade

SXPA288 User 22:04, 28 October 2016 (BST) Student 1: Matt Goodwin Student 2: Nicholas Busby Student 3: Chris Wright Student 4: Heather Allderidge

SXPA288 User 22:04, 28 October 2016 (BST) Student 2 taken control of Pirate

SXPA288 User 22:05, 28 October 2016 (BST) Conditions Cloud base: 3870m Wind speed: 29.9kmh Outside humidty: 30.5 Temp: 9.14deg C

SXPA288 User 22:14, 28 October 2016 (BST) target M39

SXPA288 User 22:17, 28 October 2016 (BST) Wind speed caused dome to shut - Peak approx 50kmh

SXPA288 User 22:30, 28 October 2016 (BST) Wind continues to peak over 50kmh. Dome remains shut

SXPA288 User 22:55, 28 October 2016 (BST) Student 3 taken control of Pirate - awaiting improvements to wind conditions

SXPA288 User 23:06, 28 October 2016 (BST)Limits adjusted to allow viewing - dome open

SXPA288 User 23:06, 28 October 2016 (BST)Outside temp 9.6C, Outside humidity 32.4%, Wind speed 37.7kph, Peak wind 44.6kph

SXPA288 User 23:06, 28 October 2016 (BST)Target M52, Filters B & G, exposure 15s achieved

SXPA288 User 23:06, 28 October 2016 (BST)Peak wind at 55kph - dome closed

SXPA288 User 23:09, 28 October 2016 (BST)Student 4 taken control of PIRATE - Queuing program ready for improvements in weather.

SXPA288 User 23:20, 28 October 2016 (BST)Outside temp - 8.95C, Outside humidity - 33.76%, Wind speed - 52.5km/h, Cloud Base - 3710m, Dome opened briefly before closing due to high wind speeds.

SXPA288 User 23:40, 28 October 2016 (BST)Dome still closed due to wind speeds peaking at over 60 km/h.

SXPA288 User 23:55, 28 October 2016 (BST)Wind speed still to high for dome to safely open. Ending session.

Richard 00:05, 29 October 2016 (BST) PIRATE set to observe science targets in the event that weather improves. Occasional bright flashes in the upper left of external camera have been seen throughout the night, attributed to lightning in the distance.

29/10/2016

Operators: Matt Pettitt, Elettra De Biasi, Anthony Palmer, David Arnot. NDA: Dean

SXPA288 User 20:12, 29 October 2016 (BST) All users able to login to PIRATE, Matt Pettitt performing the Logging Role

SXPA288 User 20:12, 29 October 2016 (BST) cloud base currently red: Fluctuating between 300-1400m

SXPA288 User 20:19, 29 October 2016 (BST) Dome still closed, students setting up schedule for M15, M39 & M52 in the meantime

SXPA288 User 20:30, 29 October 2016 (BST) Dome still closed, no change in cloud base.

SXPA288 User 20:33, 29 October 2016 (BST) Weather at 2030: Outside temp: 8.54 C, Outside Humidity: 95.43 %, Wind Speed: 26.01 km/h, Peak wind speed: 30.23 km/h, Wind direction: 167.6 degrees, Cloud base: 555.16m, Not raining

SXPA288 User 20:33, 29 October 2016 (BST) Camera temperature: -20.00 C

SXPA288 User 21:02, 29 October 2016 (BST) Dome still closed, high humidity reaching 100% and low cloud base.

SXPA288 User 21:56, 29 October 2016 (BST) Dome still closed, now raining.

SXPA288 User 22:29, 29 October 2016 (BST) Dome still closed, rain is continuing and low cloud base.

SXPA288 User 22:59, 29 October 2016 (BST) Dome still closed, rain is continuing and low cloud base.

SXPA288 User 23:30, 29 October 2016 (BST) As it is still raining, there is no longer any chance of performing any imaging this evening before midnight. Handing control back to the NDA.

30/10/2016

SXPA288 User 19:21, 30 October 2016 (GMT) Student 1 = Tom McMahon, Student 2 = Joe Webb, Student 3 - Amanda McAuley, Student 4 - Craig C

SXPA288 User 19:22, 30 October 2016 (GMT) All students logged in

SXPA288 User 19:22, 30 October 2016 (GMT) Student 3 has taken control of the observatory

SXPA288 User 19:22, 30 October 2016 (GMT) Cloud base is 3165m, which is too low for dome to open

SXPA288 User 19:30, 30 October 2016 (GMT) Student 3 creates a new template M15_3010

SXPA288 User 19:30, 30 October 2016 (GMT) waiting for cloud... cloud base is fluctuating around 3200m

SXPA288 User 19:46, 30 October 2016 (GMT) waiting for cloud, site humidity at 85.63%, this will prevent dome open even if cloud clears...

SXPA288 User 19:46, 30 October 2016 (GMT) site humidity has risen to 88%

SXPA288 User 20:20, 30 October 2016 (GMT) low cloud and high humidity continue to blight our dreams of opening the dome...

SXPA288 User 20:39, 30 October 2016 (GMT) humidity steadily climbing, now at 91.71%, the cloud base, like our spirits, is falling...

SXPA288 User 21:04, 30 October 2016 (GMT) mother nature is a cruel mistress...

SXPA288 User 21:04, 30 October 2016 (GMT) Student 3 has released control, NDA has suggested we reconvene in 1 hour.

SXPA288 User 21:04, 30 October 2016 (GMT) Student 3 is re-taking control, and switching on the light within the dome

SXPA288 User 21:04, 30 October 2016 (GMT) with the light on and outside humidity dropping by 0.10%, student excitement has been restored to previous astronomical levels.

SXPA288 User 21:05, 30 October 2016 (GMT) Student 3 has released control, we are returning at 21:30...

SXPA288 User 21:32, 30 October 2016 (GMT) and we're back sports fans, the team are ready, the cameras are cold, but the cloud is low, and the humidity is high. The dome, our portal to the heavens, remains closed.

SXPA288 User 21:32, 30 October 2016 (GMT) after a brief straw poll, we decide to return at 22:00...

SXPA288 User 22:00, 30 October 2016 (GMT) It's not looking good... We shall return again at 22:30

SXPA288 User 22:30, 30 October 2016 (GMT) The cloud base has risen substantially since 22:00, however humidity is still above operating conditions.

SXPA288 User 22:30, 30 October 2016 (GMT) We shall return at 23:00

SXPA288 User 22:52, 30 October 2016 (GMT) student 4 has taken over control

SXPA288 User 22:52, 30 October 2016 (GMT) student 4 has programmed a new template M52_20161030

SXPA288 User 22:53, 30 October 2016 (GMT) cloud has risen above the threshold, humidity is dropping, currently at 89%

SXPA288 User 23:08, 30 October 2016 (GMT) Site humidity has risen to 91.87% and cloud base has been above the threshold for 4 minutes

SXPA288 User 23:23, 30 October 2016 (GMT) Cloud base has now been ok for 10 minutes, Humidity is still to high to open the dome but has slowly been dropping

SXPA288 User 23:28, 30 October 2016 (GMT) humidity has not dropped enough for observations to be possible, we have called it a night


31/10/2016

SXPA288 User 19:30, 31 October 2016 (GMT) NDA Richard Busuttil

SXPA288 User 19:33, 31 October 2016 (GMT) Student 1 Megan McDermott Student 2 Ben Mills Student 3 Jenny Martland Student 4 Brian Maher

SXPA288 User 20:16, 31 October 2016 (GMT) All Students logged into ABOT

SXPA288 User 20:16, 31 October 2016 (GMT) Student 2 took control of ABOT dashboard

SXPA288 User 20:16, 31 October 2016 (GMT) Student 2 creates template for object M15

SXPA288 User 20:16, 31 October 2016 (GMT) Student 2 relinquished control of ABOT dashboard

SXPA288 User 20:16, 31 October 2016 (GMT) Cloud base too low for observing

SXPA288 User 20:16, 31 October 2016 (GMT) Agreed to wait break until 20:45 to see if observing conditions improve

SXPA288 User 20:45, 31 October 2016 (GMT) Cloud base at 2300 m, continues to fall

SXPA288 User 20:56, 31 October 2016 (GMT) Cloud base at 1800 m, the approaching cold front is not helping

SXPA288 User 21:00, 31 October 2016 (GMT) Group to reconvene at 21:30 for an update on conditions

SXPA288 User 21:30, 31 October 2016 (GMT) Small increase in cloud height, to a disappointing 2000 m. Giving the conditions an hour to see if the increase continues, regroup at 22:30

SXPA288 User 23:04, 31 October 2016 (GMT) Group reconvened at 22:30, cloud base rising slowly, still insufficient to begin observing. Continuing to monitor situation but it would seem unlikely to be able to obtain any images tonight

SXPA288 User 23:33, 31 October 2016 (GMT) In the absence of any rise in the cloud base the group called it a night at 23:33


01/11/2016

NDA Richard Busuttil

SXPA288 User 19:21, 1 November 2016 (GMT) Student 1 Jess, Student 2 Ben, Student 3 Max.

SXPA288 User 19:21, 1 November 2016 (GMT) All students logged into ABOT dashboard. Student 3 is taking control of ABOT dashboard.

SXPA288 User 19:21, 1 November 2016 (GMT) Dome closed, cloud base too low. Cloud base fluctuating between 2999.16 m and 3001.77 m.

SXPA288 User 20:24, 1 November 2016 (GMT) Cloud base still too low, easing time continues to reset itself to 10 minutes.

SXPA288 User 20:24, 1 November 2016 (GMT) Cloud base red, fluctuating between 3130 m and 3155.5.

SXPA288 User 20:24, 1 November 2016 (GMT) Dome opened.

SXPA288 User 20:24, 1 November 2016 (GMT) Collecting image for target M15, 30 s exposure, B and G filter, binning 2 x 2.

SXPA288 User 20:24, 1 November 2016 (GMT) Image collected for M15, 15 s exposure, B and G filter, binning 2 x 2.

SXPA288 User 20:24, 1 November 2016 (GMT) Student 2 is downloading images.

SXPA288 User 20:26, 1 November 2016 (GMT) Log logged me out and I had to copy information and log back in the Dome opened about 20:00.

SXPA288 User 20:34, 1 November 2016 (GMT) Student 1 has collected image for M15 10 s exposure, B filter, binning 2 x 2. G filter hasn't gone through.

SXPA288 User 21:10, 1 November 2016 (GMT) Cloud base 3192 dome closed, wind speed 28, Humidity 50 awaiting dome to reopen

SXPA288 User 21:18, 1 November 2016 (GMT) Student 1 scheduled template for M39, still awaiting reopening of dome.

SXPA288 User 21:33, 1 November 2016 (GMT) Dome opening, Automatically imaging schelduled template M39

SXPA288 User 21:33, 1 November 2016 (GMT) Wind speed: 26.01 Humidity: 48.97 Cloudbase 3539

SXPA288 User 21:33, 1 November 2016 (GMT) Images of M39 successfully taken, Filters: 1xG, 1xB. 30 seconds exposure

SXPA288 User 21:33, 1 November 2016 (GMT) Student 3 locating and checking M39 images

SXPA288 User 21:33, 1 November 2016 (GMT) Student 1 scheduling images for M39 filters 1xG, 1xB 15 seconds exposure

SXPA288 User 21:33, 1 November 2016 (GMT) Images successfully taken.

SXPA288 User 21:33, 1 November 2016 (GMT) Student 1 Scheduled template for M39 Filters 1xB, 1xG. 10 second exposure

SXPA288 User 21:33, 1 November 2016 (GMT) Student 3 checking quality of images in AstroImageJ

SXPA288 User 21:37, 1 November 2016 (GMT) Wind speed: 24.71 Humidity: 47.25, Cloudbase: 3567 Environment still acceptable

SXPA288 User 21:37, 1 November 2016 (GMT) Student 1 Web browser crashed, lost connection to ABOT dashboard

SXPA288 User 21:39, 1 November 2016 (GMT) Student 1 restarted browser and regained access to ABOT dashboard

SXPA288 User 21:49, 1 November 2016 (GMT) Student 3 established image quality as adequate

SXPA288 User 21:49, 1 November 2016 (GMT) Student 1 taken M39 images at 5 second exposure filters 1xB, 1xG. 2x2 binning

SXPA288 User 22:00, 1 November 2016 (GMT) Rotating roles, Environment Student 1, Student 2 Observation, Student 3 Log

SXPA288 User 22:14, 1 November 2016 (GMT) Scheduled M52 1xG 1xB 30s exposure binning 2x2

SXPA288 User 22:14, 1 November 2016 (GMT) Images taken, student 1 reviewing

SXPA288 User 22:14, 1 November 2016 (GMT) PIRATE tracking, humidity 53%, wind 27km/h, cloud base 3670m

SXPA288 User 22:19, 1 November 2016 (GMT) student 2 entered new schedule M52, 1xB 1xG, 20s exposure, 2x2 binning

SXPA288 User 22:23, 1 November 2016 (GMT) Student 1 assessing images in Astroimagej

SXPA288 User 22:29, 1 November 2016 (GMT) Student 2 PIRATE tracking, wind 42km/h, humidity 53%, cloud base 3635m

SXPA288 User 22:33, 1 November 2016 (GMT) Student 2, ordered new schedule: M15 1xB, 10s exposure, 2x2 binning

SXPA288 User 22:35, 1 November 2016 (GMT) PIRATE slewing

SXPA288 User 22:35, 1 November 2016 (GMT) Student 1 assessing images

SXPA288 User 22:44, 1 November 2016 (GMT) Student 2 released controls

SXPA288 User 22:44, 1 November 2016 (GMT) Student 3 taken control, student 1 monitoring atmosphere and log.

SXPA288 User 22:48, 1 November 2016 (GMT) Scheduled for Andromeda, ran at 20 s exposure, filters R, G, B and Luminous, 2 x 2 binning.

SXPA288 User 23:05, 1 November 2016 (GMT) Student 3 has scheduled to look at the Crab Nebula R, G, B and Luminous filters and 15 s, 2 x2 binning.

SXPA288 User 23:05, 1 November 2016 (GMT) Student 3 is now looking at Heart Nebula using R, G, B and Luminous filters with 15 s exposure, 2 x 2 binning.

SXPA288 User 23:38, 1 November 2016 (GMT) Student 3 has passed the controls over to student 2.

SXPA288 User 23:38, 1 November 2016 (GMT) Student 2 is handing over the telescope back to the NDA. Everyone has collected the data and images needed and are now signing off.

02/11/2016

DeanR 19:00, 2 November 2016 (GMT) NDA for tonight.

SXPA288 User 19:29, 2 November 2016 (GMT) Student 1 - Paul Saunders, Student 2 - Ollie Pontone, Student 3 - Salvis Gorbans

SXPA288 User 19:30, 2 November 2016 (GMT) All students logged into Observatory Dashboard

SXPA288 User 19:34, 2 November 2016 (GMT) Student 1 - Observers log, Student 2 - Taking control of ABOT, Student 3 - Weather Monitor

SXPA288 User 19:35, 2 November 2016 (GMT) Student 2 switched on the dome lights.

SXPA288 User 19:40, 2 November 2016 (GMT) Cloud cover not permitting observations.

SXPA288 User 19:46, 2 November 2016 (GMT) Average wind speed 51 km/h. Average outside temperature 9.76°C. Cloud base fluctuating between 3097 m - 3150 m.

SXPA288 User 19:53, 2 November 2016 (GMT) Cloud base improving, wind speed still too high.

SXPA288 User 20:02, 2 November 2016 (GMT) Cloud base is now 3312 m - blue light for observing. Now only waiting for wind speed to reduce.

SXPA288 User 20:10, 2 November 2016 (GMT) Cloud base is 3310 m and average wind speed is 53 km/h - red light for both.

SXPA288 User 20:23, 2 November 2016 (GMT) A ten minute tea break taken.

SXPA288 User 20:33, 2 November 2016 (GMT) Cloud base at 2995 m and average wind speed is 61 km/h.

SXPA288 User 21:02, 2 November 2016 (GMT) Average weather conditions are reducing but still not low enough for observations.

SXPA288 User 21:21, 2 November 2016 (GMT) Ten minute break taken.

SXPA288 User 21:33, 2 November 2016 (GMT) All returned from break. Observing conditions are still not favourable. Average wind speed 45 km/h. Cloud base 2749 m.

SXPA288 User 21:58, 2 November 2016 (GMT) Still no improvement on conditions. Half hour break taken.

SXPA288 User 22:32, 2 November 2016 (GMT) Back from the break. Still no improvement on conditions. Cloud base 2222 m. Average wind speed 50 km/h.

SXPA288 User 23:00, 2 November 2016 (GMT) Cloud base increased to 2485 m. Average wind remains at 50 km/h. Dome still closed and no observations able to take place.

SXPA288 User 23:14, 2 November 2016 (GMT) Wind speed reducing dramatically.

SXPA288 User 23:18, 2 November 2016 (GMT) Average wind speed clocked 45 km/h. Blue light given. Only waiting on cloud base to increase - currently at 2146 m.

SXPA288 User 23:20, 2 November 2016 (GMT) Average wind speed back to to 50 km/h. Easing time reset. Dome still closed and observations still unable to take place.

SXPA288 User 23:29, 2 November 2016 (GMT) Cloud base at an all-time low of 1938 m.

SXPA288 User 23:36, 2 November 2016 (GMT) Students have agreed to give it ten more minutes waiting time before we hand over to the NDA.

SXPA288 User 23:46, 2 November 2016 (GMT) Student 2 released control of the ABOT and Student 1 took over. Student 3 took over the log.

SXPA288 User 23:46, 2 November 2016 (GMT) Student 1 scheduled a program to observe M39. It was added to the schedule.

SXPA288 User 23:48, 2 November 2016 (GMT) Student 1 released control of the ABOT and Student 3 took control. Student 3 scheduled an observation program of M52.

SXPA288 User 23:51, 2 November 2016 (GMT) No observations able to take place in the allocated time. Observatory handed back to NDA.

SXPA288 User 23:53, 2 November 2016 (GMT) Light in dome was turned off.

03/11/2016

SXPA288 User 19:45, 3 November 2016 (GMT) Logon by Student1(Richard W), Student2(Richard H)

SXPA288 User 19:56, 3 November 2016 (GMT) Student 1(Richard H) takes control of ABOT, Student2(RW) log

SXPA288 User 20:10, 3 November 2016 (GMT) Student 2(RH) adds a new schedule to observe M15 with a G filter 30 2x2 2 and B 30 2x2

SXPA288 User 20:35, 3 November 2016 (GMT) Student 2 release control of ABOT and student 1 takes over and turns on observatory light and off and releases control of ABOT

SXPA288 User 20:56, 3 November 2016 (GMT) Sensors Wind speed, Peak wind speed, and cloud base are red

SXPA288 User 20:56, 3 November 2016 (GMT) Students logged out, observations abandoned due to bad weather

04/11/2016

NDA Stella Bradbury

Tutor 18:22, 4 November 2016 (GMT) Automated dusk flat observations were underway but aborted and dome closed in response to cloud base sensor.

SXPA288 User 19:46, 4 November 2016 (GMT) Present: Catriona Shepherd (Student 2, Log Keeper), Gurnank Cheema (Student 1, Operator), Sabine Viramoutou (Student 4, Weather Monitor), Sean Ferrin (Student 3, Data Assessor).

SXPA288 User 20:03, 4 November 2016 (GMT) Dome closed - Cloud base too low and humidity too high.

SXPA288 User 20:59, 4 November 2016 (GMT) Dome closed - Humidity at 90.93%, Cloud base at 3353.02m. Awaiting better conditions.

SXPA288 User 21:02, 4 November 2016 (GMT) Student 1 has released control.

SXPA288 User 21:04, 4 November 2016 (GMT) Student 3 has taken control.

SXPA288 User 21:05, 4 November 2016 (GMT) Dome closed. Humidity at 89.60%, Cloud base at 3457.90m. Awaiting better conditions.

'SXPA288 User 21:13, 4 November 2016 (GMT)' humidity at 73.38%; Dome still closed, cloud base at 3441.62m.

'SXPA288 User 21:15, 4 November 2016 (GMT)' Student3 has released control and pass it to Student1 ; the condition are improving.

'SXPA288 User 21:18, 4 November 2016 (GMT)' Dome is opening. conditions are reached.

'SXPA288 User 21:31, 4 November 2016 (GMT)' Student3 has submitted the program. M52_00_Filter_B_22h31.

'SXPA288 User 21:21, 4 November 2016 (GMT)' Student1 has submitted the program. M15_00_Filter_B_22h19

'SXPA288 User 21:21, 4 November 2016 (GMT)' Student1 has submitted the program. M15_00_Filter_G_22h19, Exposure times 30 seconds.

'SXPA288 User 21:21, 4 November 2016 (GMT)' Student1 has submitted the program. M15_00_Filter_Luminance_22h19, Exposure times 30 seconds.

'SXPA288 User 21:25, 4 November 2016 (GMT)' Student 2 is checking the images in Astro. The images in B & G filters look good.

'SXPA288 User 21:28, 4 November 2016 (GMT)' Student1 is releasing control & will check image quality. Student3 is pirate operator.

'SXPA288 User 21:28, 4 November 2016 (GMT)' Clouds Base is dropping again.

'SXPA288 User 21:34, 4 November 2016 (GMT)' Student3 has submitted program M52_Filter_B_22h31

'SXPA288 User 21:34, 4 November 2016 (GMT)' Student3 has submitted program M52_Filter_B_22h31. System is not responding. System is trying to focus.

'SXPA288 User 21:38, 4 November 2016 (GMT)' Wind direction is RED.Cloud Base is getting down. Dome is closing.

'SXPA288 User 21:40, 4 November 2016 (GMT)' Student3 has collected 1 image_B_Filter_ M52.

'SXPA288 User 21:49, 4 November 2016 (GMT)' Dome is opening.

'SXPA288 User 21:52, 4 November 2016 (GMT)' Student3 has submitted program M52_Filter_G_B_2images.

'SXPA288 User 22:07, 4 November 2016 (GMT)' Student3 has submitted program M39_Filter_B_22h55. 2 images. Exposure times 30 seconds.

'SXPA288 User 22:07, 4 November 2016 (GMT)' Student3 has submitted program M39_Filter_G_23h00. waiting for the images to come in Astro.

'SXPA288 User 22:07, 4 November 2016 (GMT)' Student3 received clear images for Filter_B_&_G for M52.

'SXPA288 User 22:07, 4 November 2016 (GMT)' Student3 has submitted program M39_Filter_G_clear images_ 1image.

'SXPA288 User 22:10, 4 November 2016 (GMT) Student 1 updating log. Student 2 has taken control. Weather good. Humidity 68%. submitted B&G filter for M67, Exposure times 20 seconds.

'SXPA288 User 22:20, 4 November 2016 (GMT)pirate is not responding to user command. Examining possible causes. Declination of M67 was below 30. Seeking alternative images.

'SXPA288 User 22:27, 4 November 2016 (GMT) student 2 re-imaging M52 - 2 B Filter images. Images Clear

'SXPA288 User 22:39, 4 November 2016 (GMT) Student 4 in control of telescope. Imaging M39 in B filter 30 seconds exposure. Humidity is 75%. Windspeed = 8k. Outside temp 3.22 degrees C.

'SXPA288 User 22:47, 4 November 2016 (GMT)' Images for M39 G and luminance filter images taken. All clear.

'SXPA288 User 22:52, 4 November 2016 (GMT)' Now imaging M45 in B filter, 30 second exposure. Image saturated. So attempting same image at 20 seconds.

'SXPA288 User 22:56, 4 November 2016 (GMT)' image of M45 in B filter with exposure of 10 secs and 20 secs taken. Both 10 and 20 sec exposures too saturated.

'SXPA288 User 22:59, 4 November 2016 (GMT)' image of M31, Luminance, B and G filters with 30 sec exposure. All images obtained clearly.

'SXPA288 User 23:11, 4 November 2016 (GMT)' Image of M31 with R filter, 30 second exposure. Image acceptable. Student 4 releasing control.

'SXPA288 User 23:14, 4 November 2016 (GMT) Student 2 taken control of telescope. Imaging M110, Luminance filter, 45 second exposure. Good image obtained.

'SXPA288 User 23:22, 4 November 2016 (GMT), Imaging M109, Luminance filter, 45 second exposure. Cloud base dropped below 3200 m before image could be produced.

'SXPA288 User 23:24, 4 November 2016 (GMT)' Student 2 released control of telescope.

'SXPA288 User 23:28, 4 November 2016 (GMT)' Observation session concluded.

Tutor 23:37, 4 November 2016 (GMT) Dome closed due to cloud base at the 3200 m mark. Overall star detection is 93%. Executed Waiting Transition and turned off light. Returned control to AbotVirtualObserver.

06/11/2016

Student 1 is Jessica Moule Student 2 is Alan Sully Student 3 is Lewis Saunders Student 4 is Michael Robinson

For first observation, Alan is in control of telescope, Lewis is logkeeper, Michael is doing image analysis. Jess is planned to monitor environmental conditions, but is currently having internet problems so Alan will do this.

Weather is clear, humidity at 77%

SXPA288 User 19:38, 6 November 2016 (GMT) NDA is instructing Alan in how to take first image

SXPA288 User 19:44, 6 November 2016 (GMT) Taking first text image of M15 in B and G, 30 seconds each

SXPA288 User 19:47, 6 November 2016 (GMT) Jessica emailed Lewis to say internet cable at home is damaged and we will have to proceed without her

SXPA288 User 20:00, 6 November 2016 (GMT) increased to 45 second exposure based on image analysis, taking five images of M15 in B and G

SXPA288 User 20:11, 6 November 2016 (GMT) taking 2 test images of M39 30 sec exposure, B and G

SXPA288 User 20:16, 6 November 2016 (GMT) PIRATE appears to be doing an unrecognised observing schedule, attempting to contact NDA to find out how to stop it

Site humidity above 85%, dome closing

SXPA288 User 20:19, 6 November 2016 (GMT) dome closed

SXPA288 User 20:34, 6 November 2016 (GMT) dome opening, site humidity at 81%

SXPA288 User 20:36, 6 November 2016 (GMT) observing M39, two images in each band of B and G, 30 sec

SXPA288 User 20:37, 6 November 2016 (GMT) humidity at 86%, dome closing

SXPA288 User 20:42, 6 November 2016 (GMT) switching roles: Lewis to command telescope, Alan to log, Michael on weather and image analysis

SXPA288 User 20:57, 6 November 2016 (GMT)Awaiting Dome to open at present too much humidity

SXPA288 User 21:05, 6 November 2016 (GMT) Student 3 has released scope control

SXPA288 User 21:05, 6 November 2016 (GMT) Student 3 has retaken scope control both site and outside humidity too high

SXPA288 User 21:10, 6 November 2016 (GMT) Humidity OK taking test run with M39

SXPA288 User 21:25, 6 November 2016 (GMT) Overexposed reducing time to 10 sec

SXPA288 User 21:36, 6 November 2016 (GMT)Humidity High dome now closing again

SXPA288 User 22:06, 6 November 2016 (GMT) Humidity climbing still

SXPA288 User 22:32, 6 November 2016 (GMT) Outside humidity starting to reduce but still above 80%

SXPA288 User 22:32, 6 November 2016 (GMT) Student 3 releasing control of scope

SXPA288 User 22:49, 6 November 2016 (GMT) Student 4 has taken control

SXPA288 User 22:55, 6 November 2016 (GMT) outside humidity shows blue but still over 80% contacting NDA

SXPA288 User 23:20, 6 November 2016 (GMT) Dome open taking tests off M39

SXPA288 User 23:22, 6 November 2016 (GMT) Tanking 5 in both filters at 5 secs

SXPA288 User 23:22, 6 November 2016 (GMT) Still too exposed taking 5 at 1 Sec

SXPA288 User 23:28, 6 November 2016 (GMT) Now trying M52 at 30 secs

SXPA288 User 23:28, 6 November 2016 (GMT) Doing M52 X 5 at 30 Secs

SXPA288 User 23:48, 6 November 2016 (GMT) Cloud base and humidity not correct dome closed

SXPA288 User 23:48, 6 November 2016 (GMT) Handing back telecescope


07/11/2016

NDA is Stella Bradbury

Tutor 18:59, 7 November 2016 (GMT) Bias and Dark Runs were taken but dome remains closed as the telescope is very clearly IN cloud on the external camera and registering rain.

SXPA288 User 19:40, 7 November 2016 (GMT) Observing Team for the evening: Student 1: Jessica Stoneman (expected around 9pm), Student 2: Carleen Duffy, Student 3: Alvin Maraj, Student 4: David Hewitt(Log editor)

SXPA288 User 19:54, 7 November 2016 (GMT) Dome closed due to cloud base (approximately 1.05m), site humidity (approximately 100.00%) and rain.

SXPA288 User 20:04, 7 November 2016 (GMT) All star camera detects 1 star in the west

SXPA288 User 20:04, 7 November 2016 (GMT) Dome closed due to cloud base (approximately 0.00m), site humidity (approximately 100.00%) and rain.

SXPA288 User 20:10, 7 November 2016 (GMT) Student 2 has taken control and has switched on the IR camera.

SXPA288 User 20:35, 7 November 2016 (GMT) Student 2 has added a simple schedule for M39 at 30 and 45 second intervals in B and G.

SXPA288 User 20:52, 7 November 2016 (GMT) Student 2 has released control of the telescope

SXPA288 User 20:52, 7 November 2016 (GMT) Student 2 has taken over as log editor.

SXPA288 User 20:53, 7 November 2016 (GMT) Student 4 has taken control of the telescope.

SXPA288 User 20:53, 7 November 2016 (GMT) Dome closed due to Cloud base (1.05m), humidity (100%) and rain.

SXPA288 User 20:55, 7 November 2016 (GMT) Student 4 has added a simple schedule for M15 at 45 and 60 second intervals in B and G.

SXPA288 User 21:01, 7 November 2016 (GMT) Student 4 has released control of the telescope.

SXPA288 User 21:05, 7 November 2016 (GMT) Student 3 has taken control of the telescope.

SXPA288 User 21:15, 7 November 2016 (GMT) Student 3 has added a simple schedule for M29 at 30 second intervals in B and G.

SXPA288 User 21:30, 7 November 2016 (GMT) Student 3 has switched off the IR light.

SXPA288 User 21:40, 7 November 2016 (GMT) Student 3 has taken over as log editor.

SXPA288 User 21:43, 7 November 2016 (GMT) there is no improvement in environmental conditions, cloud base (approximately 1.05m), site humidity (approximately 100.00%). NDA has advised that weather condition will likely not improve. Lab ended

Tutor 22:17, 7 November 2016 (GMT) Returned control to AbotVirtualObserver. IR is off, camera is at -20C, dome is closed. WASP follow-up observations are queued but telescope is still in cloud. Can almost smell the damp bushes from here.

11/11/2016

Additional SXPA session without NDA support

Ulrich Cloud base lifted at around 22 GMT, wind stopped observations at about 23:30pm. SXPA288 student 4 obtained data for M36, M37

12/11/2016

Additional SXPA session without NDA support

SXPA288 User 20:03, 12 November 2016 (GMT) Sky clear - setting up open clusters

SXPA288 User 20:06, 12 November 2016 (GMT) [Alan Cayless, logged in as SXPA288.ModChair] M29 - 4,8, 15, 30 Second exposures

SXPA288 User 20:16, 12 November 2016 (GMT) Logged out, logged back in again as Student 4

SXPA288 User 20:22, 12 November 2016 (GMT) M15 - 15, 30, 60, 120 second exposures

SXPA288 User 21:07, 12 November 2016 (GMT) M2 - 15, 30, 60, 120 second exposures

SXPA288 User 21:22, 12 November 2016 (GMT) Moon (1 day before full) coming into view on all-sky camera

SXPA288 User 21:48, 12 November 2016 (GMT) M52 HDR - 1, 10, 100 Second exposures

SXPA288 User 22:04, 12 November 2016 (GMT) M39 HDR - 1, 10, 100 Second exposures

SXPA288 User 22:38, 12 November 2016 (GMT) M29 HDR - 1, 10, 100 Second exposures

SXPA288 User 22:45, 12 November 2016 (GMT) Observing run finished. Release control. Signed out.

14/11/2016

Supplementary SXPA288 session

NDA on-line is Stella Bradbury

Tutor 18:59, 14 November 2016 (GMT) Observatory currently in cloud - rain sensor triggered.

SXPA288 User 19:20, 14 November 2016 (GMT)Student 1 (Craig Cant) and Student 3 (Amanda McAuley) have logged onto PIRATE dashboard

SXPA288 User 19:26, 14 November 2016 (GMT)It's raining and humidity is around 96% with 0m cloud cover. Taking break till 20.00

SXPA288 User 20:08, 14 November 2016 (GMT)Still raining. We'll come back at 21.00

SXPA288 User 21:07, 14 November 2016 (GMT)Still raining. We'll try again at 22.00

SXPA288 User 22:25, 14 November 2016 (GMT)Rain has stopped and humidity is falling, but cloud cover is still very low (approx 18m) and doesn't seem to be rising soon. 9% of stars detected in all sky camera. Signing off. No observations taken.

Tutor 22:33, 14 November 2016 (GMT) Control returned to Virtual Observer. WASP targets queued. Awaiting shift in cloud base currently at 26 m (and clearly visible in bright moonlight).

03/03/2017

S382 Alpheratz Group
Observers: Gavin Duck; Darren Smith; Lester Sandles; Jonathan Ritchie
PIRATE Operator: Darren Smith; Log Keeper: Gavin Duck

Alpheratz 20:33, 3 March 2017 (GMT) All environmental factors acceptable. Outside Temp: 3.56C. Outside humidity: 8.83%. Average wind speed 20.32km/h (peak: 28.97 km/h). Cloud base: 4160.37 m. Not raining. Dome temperature 1: 2.74C; 2: 3.38C. Dome humidity 1: 24.24%; 2: 22.87%. Dome dew point 1: -15.57C; 2: -15.77C. Site brightness: 0.12 W/m^2. Site dust level: 0.00050 m^-3. [entry updated Alpheratz 21:09, 3 March 2017 (GMT)]

Alpheratz 20:48, 3 March 2017 (GMT) Test observations on target J052705.88+213525.5. RGB filters tested with 2 exposures each at 20s.

Alpheratz 20:59, 3 March 2017 (GMT) Test schedule submission not appearing on queue. NDA exploring problems.

Alpheratz 21:15, 3 March 2017 (GMT) NDA (Meredith Morrell) took control of PIRATE at c.2058 to attempt to rectify problem. Released control at this time stamp.

Alpheratz 21:23, 3 March 2017 (GMT) Control retaken. Problem judged to be test target too close to Moon.

Alpheratz 21:26, 3 March 2017 (GMT) Test schedule edited to actual target KM UMa (J114749.04+351335.2) as altitude now above 30 degrees. Two exposures each of RGB filters at 20s each.

Alpheratz 21:27, 3 March 2017 (GMT) Schedule saved not submitted. Control relinquished to allow software engineer to take control.

Alpheratz 22:05, 3 March 2017 (GMT) At 21:49, observations began.

Alpheratz 22:07, 3 March 2017 (GMT) Control regained. Problem appears to be a software issue concerning time schedule was submitted compared to saved. Work around to define a timed observation schedule.

Alpheratz 22:21, 3 March 2017 (GMT) Observations halted at 2204.

Alpheratz 22:24, 3 March 2017 (GMT) 36 images sent to AstroDrive under Alpheratz_Test_Exposure.

Alpheratz 22:24, 3 March 2017 (GMT) Counts: R ~ 32,000; B ~ 15,000; G ~ 22,000.

Editor of Observer Log changed to Darren Smith. Operator changed to Lester Sandles

Alpheratz 22:36, 3 March 2017 (GMT) Our aim is to achieve exposure times for each filter that gives maximum exposure values of 10,000 each.

Alpheratz 22:41, 3 March 2017 (GMT) For our second test run we are setting exposure times of R-7s G-10s B-14s

Alpheratz 22:47, 3 March 2017 (GMT) A mistake with the co-ordinates in Test Exposure 2 was spotted so run was terminated.

Alpheratz 22:50, 3 March 2017 (GMT) Test Exposure 3 was started with correct co-ordinates for KM UMa

Alpheratz 22:58, 3 March 2017 (GMT) Test Exposure 3 gave counts of R ~ 7500 G ~ 8700 B ~ 8000.

Alpheratz 22:58, 3 March 2017 (GMT) From these results we adjusted the exposure times of Test Exposure 4 to R- 9s G- 11s B- 18s

Due to technical problems we were unable to use the observer's log for the rest of the evening. Records were kept offline which are presented below

Alpheratz 18:01, 19 March 2017 (GMT) The values for for test 4 were R ~ 6700 G ~ 6500 B ~ 7300. Since these values were surprisingly low we run the test again in case we had got unlucky with a stray cloud.

Alpheratz 18:01, 19 March 2017 (GMT) For Test Exposure 5 the values are R ~ 10800 G ~ 9800 B ~ 12000. The result for Red and Green are fine but a bit high for Blue

Alpheratz 18:01, 19 March 2017 (GMT) Test 6 repeats the values for Red And Green to check they are OK and Blue will be reduced to 16s

Alpheratz 18:01, 19 March 2017 (GMT) The test 6 values were R ~ 12000 G ~ 14000 B ~ 14000

Alpheratz 18:01, 19 March 2017 (GMT) These values were used for the final run that ran all night

04/03/2017

Vega 19:31, 4 March 2017 (GMT)
Nick Evans; Giles Charter; Sam Arnsby; Paul Moore

Vega 19:40, 4 March 2017 (GMT) Environmental conditions acceptable for observing. Not raining. Outside temperature: ~ 4 degrees. Outside humidity: ~ 69%. Average wind speed: ~ 20km/h (peak speed: ~ 25km/h). Cloud base: ~ 3300m.

Vega 20:17, 4 March 2017 (GMT) Acquired test source for calibrating exposure times. SWASP J051740.27+335355.4 coordinates RA 5 17 40.27 / DEC 33 53 55.39

Vega 20:19, 4 March 2017 (GMT) Giles Charter telescope operator. Nick Evans log keeper. Sam Arnsby and Paul Moore checking image frame quality and environmental conditions.

Vega 20:20, 4 March 2017 (GMT) Set Red filter 10s exposure; Green filter 10s exposure; Blue filter 10s exposure

Vega 20:41, 4 March 2017 (GMT) Checked image frames. All underexposed. Pixel counts: red 5258, green 3315, blue 1100

Vega 20:43, 4 March 2017 (GMT) Set Red filter 15s exposure; Green filter 15s exposure; Blue filter 15s exposure

Vega 20:49, 4 March 2017 (GMT) Checked image frames. All underexposed. Pixel counts: red 7998, green 5295, blue 3900

Vega 20:49, 4 March 2017 (GMT) Set Red filter 20s exposure; Green filter 25s exposure; Blue filter 30s exposure

Vega 21:10, 4 March 2017 (GMT) Checked image frames. All underexposed.

Vega 21:10, 4 March 2017 (GMT) Set Red filter 25s exposure; Green filter 30s exposure; Blue filter 35s exposure

Vega 21:10, 4 March 2017 (GMT) Checked image frames. All underexposed.

Vega 21:16, 4 March 2017 (GMT) Set Red filter 30s exposure; Green filter 35s exposure; Blue filter 40s exposure

Vega 21:28, 4 March 2017 (GMT) Checked image frames. All underexposed. Pixel counts: red 9500, green 11000, blue 10500

Vega 21:28, 4 March 2017 (GMT) Set Red filter 40s exposure; Green filter 45s exposure; Blue filter 50s exposure

Vega 21:28, 4 March 2017 (GMT) Environmental conditions stable

Vega 21:36, 4 March 2017 (GMT) Checked image frames. Good exposures. 10000-14000 pixel counts for all of red, green, blue. Agreed to test increasing exposures times by a further 10 seconds for each filter to aim for 15000-18000 pixel counts.

Vega 21:36, 4 March 2017 (GMT) Set Red filter 50s exposure; Green filter 55s exposure; Blue filter 60s exposure

Vega 21:46, 4 March 2017 (GMT) Checked image frames. Good exposures. Pixel counts: red 15945, green 17778, blue 13636. Decided to increase red by a further 2s and blue by 10s, keep green at 55s

Vega 21:46, 4 March 2017 (GMT) Set Red filter 52s exposure; Green filter 55s exposure; Blue filter 70s exposure

Vega 21:54, 4 March 2017 (GMT) Checked image frames. Good exposures. Pixel counts: red 17956, green 18703, blue 16793. Decided to increase blue by a further 2s, keep red at 52s and green at 55s

Vega 21:54, 4 March 2017 (GMT) Set Blue filter 72s exposure

Vega 22:00, 4 March 2017 (GMT) Checked image frames. Pixel count: blue 15899

Vega 22:00, 4 March 2017 (GMT) Set Blue filter 80s exposure

Vega 22:14, 4 March 2017 (GMT) OK with all test exposures based on pixel counts. Red 52s, Green 55s, Blue 80s. Pixel counts all in the 16000-18000 range.

Vega 22:25, 4 March 2017 (GMT) Acquired target. SWASP J132837.25+353311.6 coordinates RA 13 28 37.25 / DEC 35 33 11.59. Red 52s, Green 55s, Blue 80s. 2x2 binning. 3 repeats.

Vega 22:35, 4 March 2017 (GMT) Sequence set for observations throughout night

Vega 22:46, 4 March 2017 (GMT) Check of first few images to arrive in each filter. All in the 7000-8500 pixel count range with target at 30 degrees altitude. Images confirmed as within good pixel count range given the altitude so as not to be saturated when star at the zenith. Sequence left to run.

06/03/2017

Capella team John Deacon Operator, Matt Carden Data checker and analyst, Daniel Chiswell Environmental and weather, Howard Mitchell Log keeper.

Capella 19:32, 6 March 2017 (GMT) Dome closing due cloud Capella 19:33, 6 March 2017 (GMT) Joe McKay joins

Capella 19:44, 6 March 2017 (GMT) Outside temp 9.92C, Humidity 23%, Avg Wind speed 17kmh, Peak wind speed 22kmh, Wind direction 85 deg, Avg wind dir 89, cloud base 3500m, no rain, dome temp 11.4 C dome humidity 32.5%, dome dewpoint -4.4C

Capella 19:45, 6 March 2017 (GMT) Ulrich and Fiona briefing call starts

Capella 20:05, 6 March 2017 (GMT) Ulrich outlined where to find the calibration frames in AstroDrive

Capella 20:11, 6 March 2017 (GMT) Dome opens

Capella 20:26, 6 March 2017 (GMT) Camera at -20C

Capella 20:26, 6 March 2017 (GMT) Schedule system has been updated, so need to check that jobs can be submitted

Capella 20:26, 6 March 2017 (GMT) conditions stable no change. No rain

Capella 20:26, 6 March 2017 (GMT) Adding simple schedule for exposure trial target J053807.15+422029.1

Capella 20:26, 6 March 2017 (GMT) RA 05 38 07.15 Dec +42 20 29.1

Capella 20:26, 6 March 2017 (GMT) Dome closed

Capella 20:26, 6 March 2017 (GMT) Preparing schedules

Capella 20:52, 6 March 2017 (GMT) Capella20:45, 6 March 2017 (GMT) Role change

McKay: Log keeper, Matt: Web cams, Howard: Analyst, John: Weather monitor, Daniel: Operator.

Capella 20:52, 6 March 2017 (GMT)McKay recording.

Capella20:50, 6 March 2017 (GMT) Group talking through scheduling procedure.

Capella 20:52, 6 March 2017 (GMT)Schedule templates being organised.

Capella 20:54, 6 March 2017 (GMT)Weather and sensors report - all within limits.

Capella 20:54, 6 March 2017 (GMT)Schedule and Filters discussion

Capella 20:58, 6 March 2017 (GMT)Clear sky seen overhead.

Capella 21:02, 6 March 2017 (GMT)Cloud base has turned Blue. Dome opening.

Capella 21:02, 6 March 2017 (GMT)Queue cleared

Capella 21:05, 6 March 2017 (GMT)Ulrich joined group

Capella 21:07, 6 March 2017 (GMT) Schedules commencing for 1SWASP J05307.5 +422029.1 (Calibration Star)

Capella 21:10, 6 March 2017 (GMT) Group organising image parameters

Capella 21:13, 6 March 2017 (GMT)Ulrich left group. Back in 30 mins

Capella 21:19, 6 March 2017 (GMT) Group having trouble downloading images from source

Capella 21:49, 6 March 2017 (GMT)Role Change
Matt: Log Keeper, John: Webcams, Dan: Analyst, John: Weather Monitor, Howard: Operator, Joe McKay: Suspected technical issues

Capella 21:49, 6 March 2017 (GMT)Weather and sensors report - all within limit

Capella 21:49, 6 March 2017 (GMT)Calibration star not easy to discern in images, changed target to 1SWASP J112541.63+423449.6

Capella 21:53, 6 March 2017 (GMT)Set exposure to R 40s, R80s, R160s Capella 22:23, 6 March 2017 (GMT)40s filter looks about right but may be a bit too bright at culmination - try R 30s, G 30&40s, B 30&40s

John: Log keeper, Howard: Monitor, Matt: Operator, Daniel: Analyst

Capella 22:42, 6 March 2017 (GMT) R: 30s gives 12250 pixels Capella 22:42, 6 March 2017 (GMT) Weather good - cloud cover 3700 m Capella 22:53, 6 March 2017 (GMT) G: 40s gives 13442 pixels, B: indeterminate - run again Capella 22:53, 6 March 2017 (GMT) Run B again at 80s, 120s and 140s Capella 22:53, 6 March 2017 (GMT) Dome closed - cloud cover - base 2960 m Capella 23:24, 6 March 2017 (GMT)Will go for R: 30s, G: 40s B: 100s Capella 23:24, 6 March 2017 (GMT) NDA intervention - Telescope is taking images with the dome closed. Capella 23:24, 6 March 2017 (GMT) Telescope restarted Capella 23:24, 6 March 2017 (GMT) Cloud base down to 1700 m Capella 23:58, 6 March 2017 (GMT) Problem with weather sensor connections - may not be able to run even if the weather improves and the dome opens again. Capella 23:58, 6 March 2017 (GMT) Cloud base still 1900 m. Set the schedule on a timed run with the above exposures with a repeating 3 x red, 3 x green, 3 x blue cycle. Capella 23:58, 6 March 2017 (GMT) Cleared queue and submitted.


07/03/2017

S382 PIRATE Project Deneb Group
Users : Peter Aston, Andrew Hunt, David Greenway, Maureen Nordon 
Deneb 20:59, 7 March 2017 (GMT) Main Target Star : J114749.04+35 13 35.2

Deneb 20:11, 7 March 2017 (GMT) logger : Maureen

Deneb 20:11, 7 March 2017 (GMT) Telescope Operator : Peter

Deneb 20:11, 7 March 2017 (GMT) Data analyst: Andy

Deneb 20:11, 7 March 2017 (GMT) Environmental Sensors : David

Deneb 20:19, 7 March 2017 (GMT) Scheduling templates for test star 6h 33.6, + 35 21 17.3, exposure time : 20s, background 1400 R filter

Deneb 20:19, 7 March 2017 (GMT) st exposure too dark but adjusting contrast, possibly acceptable

Deneb 20:36, 7 March 2017 (GMT) Image acceptable only 1 pixel saturated but at far end of field

Deneb 20:36, 7 March 2017 (GMT) B filter 30s

Deneb 20:36, 7 March 2017 (GMT) Test star not centralised. Educated guess.

Deneb 20:36, 7 March 2017 (GMT) All weather parameters good. Very light winds and clear skies

Deneb 20:36, 7 March 2017 (GMT) Completed imaging for B filter

Deneb 20:36, 7 March 2017 (GMT) Indicating 6 templates completed

Deneb 20:59, 7 March 2017 (GMT) Repeating schedule with B filter

Deneb 20:59, 7 March 2017 (GMT) Images coming through

Deneb 20:59, 7 March 2017 (GMT) Scheduling exposures R 30s, B 40s.

Deneb 20:59, 7 March 2017 (GMT) Image Peak 30000

Deneb 20:59, 7 March 2017 (GMT) Image peak B 40s 12000 Dimmer stars 6000

Deneb 20:59, 7 March 2017 (GMT) Image peak R 30s 6000-10000 Dimmer stars- 2-3000

Deneb 21:05, 7 March 2017 (GMT) Image peak B 50s, 19000 , Dimmer stars 4000

Deneb R 40s target area 16000, Dimmest stars 4000. G 45s 20000 Dimmest star 2500

Deneb 21:19, 7 March 2017 (GMT) David scheduled new template for Target Star J11 47 49.04 +35 13 35.2. 2x2 binning, repeat x2, B 50s, R 40s, G 45s.

Deneb 21:44, 7 March 2017 (GMT) Checking drive- Andy

Deneb 21:44, 7 March 2017 (GMT) !st Image through

Deneb 21:44, 7 March 2017 (GMT) Peak 21-30000 at stars of interest in B 50s

Deneb 21:44, 7 March 2017 (GMT) Peak 40000 at centre stars down to 4000 G 45s

Deneb 21:44, 7 March 2017 (GMT) Peak 43000 at centre stars down to 4000 R 40s

Deneb 21:44, 7 March 2017 (GMT) New schedule: Peter .B 50s, G 45s, R 40s. 2x2 binning repeat x 5. Time based.

Deneb 21:44, 7 March 2017 (GMT) Submitted

Deneb 21:48, 7 March 2017 (GMT) data coming through

Deneb 21:48, 7 March 2017 (GMT) Weather stable

Deneb 21:56, 7 March 2017 (GMT) Images good

Deneb 22:04, 7 March 2017 (GMT) Peter, David logged off. Andy monitoring

Deneb 22:07, 7 March 2017 (GMT) Maureen logging off

Deneb 14:16, 9 March 2017 (GMT) The following entries were transferred on 9 March from paper log.

Deneb 22:49, 7 March 2017 (GMT) Detecting possible saturation problems. Target star close to limit and others at 40000

Deneb 23:02, 7 March 2017 (GMT) NDA intervenes re zenith time.Peak value ( target star ) needs to be <45k

Deneb 23:18, 7 March 2017 (GMT) May need to drop exposure times

Deneb 23:20, 7 March 2017 (GMT) on G we have 57000

Deneb 23:24, 7 March 2017 (GMT) NDA double checked that we were measuring target star

Deneb 23:35, 7 March 2017 (GMT) New schedule submitted: B 40s , G 35s , R 30s. All x5

Deneb 23:38, 7 March 2017 (GMT) Getting 49k as a peak. decision to drop further

Deneb 23:46, 7 March 2017 (GMT) NDA said if earlier images had target star < 55k. Images still usable

Deneb 23:46, 7 March 2017 (GMT) G at 22k

Deneb 23:48, 7 March 2017 (GMT) 23.48 B at 28k

Deneb 00:05, 8 March 2017 (GMT) 00.05 B at 23k , G at 30k

Deneb 00:09, 8 March 2017 (GMT) 00.09 Session ends confident of securing usable data. Andy to monitor for hour.


08/03/2017

Observing Team: Simon Bell, Andrew Bowers, Adam Nawal, Joe Spencer and Sami Zenasni. Pollux 19:35, 8 March 2017 (GMT) Logger: Simon Bell

Pollux 19:37, 8 March 2017 (GMT) Operator: Adam Nawal

Pollux 19:40, 8 March 2017 (GMT) Environmental Sensors: Sami Zenasni

Pollux 19:52, 8 March 2017 (GMT) Data: Andy Bowers, Joe Spencer.

Pollux 19:54, 8 March 2017 (GMT) All users connected to Abot.

Pollux 19:58, 8 March 2017 (GMT) Ulrich joined Skype group chat for briefing.

Pollux 20:00, 8 March 2017 (GMT) Dome opened.

Pollux 20:08, 8 March 2017 (GMT) Tested a different target star to help determine exposure time (coordinates: 05 38 07.15 +42 20 29.1).

Pollux 20:17, 8 March 2017 (GMT) Outside temp 9.97ºC, Outside Humidity 21.29%, Avg Wind speed 16.02km/h, Peak wind speed 19.35km/h, Wind direction 120.70º, Avg wind dir 119.02º, cloud base 3756.18m, no rain, dome temp 9.54ºC dome humidity 31.09%, dome dewpoint -6.70ºC.

Pollux 20:19, 8 March 2017 (GMT) Initial schedule submitted.

Pollux 20:25, 8 March 2017 (GMT) For Pollux_Initial, exposure appears too bright and the picture is very 'busy'. Finder charts suggested to fix the issue.

Pollux 20:32, 8 March 2017 (GMT) Ulrich left discussion, NDA on standby.

Pollux 20:35, 8 March 2017 (GMT) For test star, exposure currently set at 20s for each filter.

Pollux 21:07, 8 March 2017 (GMT) Exposure altered: set to 30s for each filter. Target peak count rate is in the 30000-40000 range.

Pollux 21:12, 8 March 2017 (GMT) Images received for all 3 filters (blue, red, green) - peak count now at approximately 15000 for test star.

Pollux 21:15, 8 March 2017 (GMT) Finder chart for test star used through STScI website. Team working on adjusting the exposure time.

Pollux 21:35, 8 March 2017 (GMT) At exposure time of 120s, G:38000, R:54000, B:25000

Pollux 22:33, 8 March 2017 (GMT) G: 100s gives 28000, R: 90s gives 36000, B: 120s gives 19000. Agreed exposure times for the 3 filters.

Pollux 22:54, 8 March 2017 (GMT) Outside temp 9.53ºC, Outside Humidity 28.03%, Avg Wind speed 27.35km/h, Peak wind speed 31.62km/h, Wind direction 106.01º, Avg wind dir 106.09º, cloud base 3787.62m, no rain, dome temp 8.88ºC, dome humidity 36.08%, dome dewpoint -5.31ºC.

Pollux 23:24, 8 March 2017 (GMT) Consulted NDA - Set schedule on a timed run with the exposure time stated above.

Pollux 23:27, 8 March 2017 (GMT) Weather conditions are good.

Pollux 23:35, 8 March 2017 (GMT) All 3 filters giving count rates between 16000-20000.

Pollux 23:36, 8 March 2017 (GMT) Observing team signed off.


09/03/2017

Target Star: 1SWASP J112541.63 +423449.6

Arcturus 20:03, 09 March 2017 (BST) Observers: Alex Roberts, Andy Bettger, Daniel Toppin, Alex Davidson

PIRATE operator: Alex Davidson; Log Keeper: Daniel Toppin
Exposure Time: 20 s
Filter: Red, Green, Blue

Arcturus 19:56, 09 March 2017 (BST) Dome Opened, Humidity 29%

Arcturus 20:32, 09 March 2017 (BST) Entering schedule for target

Arcturus 21:43, 9 March 2017 (GMT)

Arcturus 20:37, 09 March 2017 (BST) Alex D Taking control of PIRATE

Arcturus 20:37, 09 March 2017 (BST) Alex D clearing queue

Arcturus 20:40, 09 March 2017 (BST) Telescope tracking

Arcturus 20:43, 09 March 2017 (BST) Observing

Arcturus 20:50, 09 March 2017 (BST) Issues downloading images from Astro Drive

Arcturus 20:53, 09 March 2017 (BST) Renaming template name 2nd request and submitted

Arcturus  20:55 March 2017 (BST) Decreased exposure time to 30s

Arcturus 21:01March 2017 (BST) Peak is now 4000 on blue filter

Arcturus 21:01 March 2017 (BST) Peak is now 6700 on green filter

Arcturus 21:01 March 2017 (BST) Peak is now 5200 on red filter

Arcturus 21:05 March 2017 (BST) third request submitted with 60s exposure time

Arcturus 21:05 March 2017 (BST) RGB filters

Arcturus 21:08 March 2017 (BST) First image for the third request received

Arcturus 21:08 March 2017 (BST) Peak is now 16000 on red filter

Arcturus 21:11 March 2017 (BST) Peak is now 17200 on green filter

Arcturus 21:08 March 2017 (BST) Peak is now 6700 on blue filter

Arcturus 21:20 March 2017 (BST) Changing naming convention for files

Arcturus 21:25 March 2017 (BST) R filter peak is still at 16000

Arcturus 21:25 March 2017 (BST) G filter peak still at 17200

Arcturus 21:25 March 2017 (BST) Awaiting blue filter results at 120s

Arcturus 21:08 March 2017 (BST) B filter is at 13146 at exposure of 120s

Arcturus 21:48, 9 March 2017 (GMT)Contact NDA to confirm long run end time

Arcturus 21:55, 9 March 2017 (GMT) Long run scheduled file named Long_Run_Schedule_Arcturus_090317

Arcturus 21:55, 9 March 2017 (GMT) Filters set for R 60, G 60, B 120

Arcturus 21:58, 9 March 2017 (GMT) Long Schedule set

Arcturus 22:11, 9 March 2017 (GMT) Peak values reducing

Arcturus 22:21, 9 March 2017 (GMT) High cloud visible on all sky cam

Arcturus 22:44, 9 March 2017 (GMT) B filter peak is at 22973

Arcturus 22:44, 9 March 2017 (GMT) R filter peak is at 28705

Arcturus 22:44, 9 March 2017 (GMT) G Filter peak is at 17973

Arcturus 22:46, 9 March 2017 (GMT) Target is currently at 55 degrees

Arcturus 23:03, 9 March 2017 (GMT) Sensor Data Recorded Outside Temp: 9.14 deg Outside Humidity: 30.00% Wind Speed:28.97km/h Average Wind speed: 30.21km/h Peak Wind Wpeed: 38.81km/h Wind Direction: 134.15 deg Average Wind Direction: 133.77deg Cloud Base: 3588.43m Rain: Not raining Weather Device Temp: 8.24 deg Dome Temp 1: 8.90 deg Dome Humidity 1: 37.43% Dome Dew Point 1: -4.81 deg Dome Temp 2: 8.34 deg Dome Humidity 2: 37.24% Dome Dew Point 2: -5.40 deg Light Level: 0.00 w/m^2

Arcturus 23:14, 9 March 2017 (GMT) All counts between 17500 -19500

Arcturus 23:46, 9 March 2017 (GMT) B filter peak is at 12700

Arcturus 23:46, 9 March 2017 (GMT) G Filter peak is at 12700

Arcturus 23:46, 9 March 2017 (GMT) R Filter peak is at 14700

10/03/2017

Observers : Matt Carden, Joe English, Howard Mitchell, Anne-Marie Dunbar, Nicola McIntyre

Capella 19:57, 10 March 2017 (GMT)Nicola : observers log

CapellaJoe : operator and analysis

Capella 19:57, 10 March 2017 (GMT)Anne-Marie : sensors

Capella 19:57, 10 March 2017 (GMT)Matt : background analysis

Capella 19:57, 10 March 2017 (GMT)Howard : background analysis

Test star : J053807.15 +42 20 29.1

Capelladome closed

Capella 19:57, 10 March 2017 (GMT)sensor check : camera (-20) outside temp (9.73) outside humidity (25.84) wind speed (11.26 km p h) cloud base (3593.62 m) no rain

Capella 19:57, 10 March 2017 (GMT)dome opened

Capellasetting test images

Capella 19:59, 10 March 2017 (GMT)R filter 30 s 2 x 2 binning

Capella 19:59, 10 March 2017 (GMT)G filter 40 s 2 x 2 binning

Capella 19:59, 10 March 2017 (GMT)B filter 100 s 2 x 2 binning

Capella 20:04, 10 March 2017 (GMT)setting up test run for target star J112541.63+423449.6

Capella 20:04, 10 March 2017 (GMT)R filter 30 s 2 x 2 binning

Capella 20:04, 10 March 2017 (GMT)G filter 40 s 2 x 2 binning

Capella 20:04, 10 March 2017 (GMT)B filter 100 s 2 x 2 binning

Capella 20:11, 10 March 2017 (GMT)checked IR off

Capella 20:11, 10 March 2017 (GMT)cleared queue

Capella 20:11, 10 March 2017 (GMT)submitted test runs

Capella 20:26, 10 March 2017 (GMT)sensor check : outside temp (9.76) outside humidity (25.6%) wind speed (12.09 km p h) cloud base (3596) no rain

Capella 20:26, 10 March 2017 (GMT)Joe checking images for test star

Capella 20:26, 10 March 2017 (GMT)retrying R and G filters to 50 s and B filter staying at 100 s

Capella 20:26, 10 March 2017 (GMT)second test run submitted

Capella 20:26, 10 March 2017 (GMT)sensor check : outside temp 9.6 humidity 25.68% wind speed 13.23 km p h cloud base 3520.25 no rain

Capella 20:36, 10 March 2017 (GMT)Joe checking test images

Capella 20:36, 10 March 2017 (GMT)retrying R and G filters to 50 s and B filter staying at 100 s for 2nd run

Capella 20:43, 10 March 2017 (GMT)2nd run for test star submitted

Capella 20:47, 10 March 2017 (GMT)sensor check : outside temp 9.58 outside humidity 25.53% wind speed 11.96 km p h cloud base 3541.25 m no rain

Capella 20:47, 10 March 2017 (GMT)target star now over 30 degree mark

target star : J112541.63+423449.6

Capella 20:48, 10 March 2017 (GMT)Joe submitting test run for target star R and B filters 50 s, B filter 100 s

Capella 20:56, 10 March 2017 (GMT)inspecting test images

Capella 21:11, 10 March 2017 (GMT) Role Change

Operator/analysis: Anne-Marie Dunbar, Log Keeper: Joe English, Weather/Webcams: Nicola McIntyre, Analysts: Howard Mitchell, Matt Carden

Capella 21:23, 10 March 2017 (GMT) Sensor Check: Outside Temp 9.29; Outside Humidity 24.90%, Wind Speed 14 kmph, Cloud Base 3344.10 m, No Rain

Capella 21:24, 10 March 2017 (GMT) Preparing Time-Based Schedule for Target Star

Capella 21:28, 10 March 2017 (GMT) Schedule prepared with the following exposure times: Red - 50s, Green - 50s, Blue - 100s

Capella 21:31, 10 March 2017 (GMT) Counts on initial test on target using same parameters: R - 20,101; G - 15,149; B - 15,563

Capella 21:32, 10 March 2017 (GMT) Time-based schedule to run from 20:00 to 06:25. Consultation with NDA. Schedule approved.

Capella 21:39, 10 March 2017 (GMT) Monitoring the initial frames being recorded

Capella 21:50, 10 March 2017 (GMT) Sensor Check: Outside Temp 9.4; Outside Humidity 22%, Wind Speed 17 kmph, Cloud Base 3577 m, No Rain

Capella 22:15, 10 March 2017 (GMT) Role Change

Operator/analysis: Nicola McIntyre, Log Keeper: Anne-Marie Dunbar, Weather/Webcams: Joe English, Analysts: Howard Mitchell, Matt Carden

Capella 22:21, 10 March 2017 (GMT) Sensor Check: Outside Temp 9.7; Outside Humidity 22%, Wind speed 11.8 kmph, Cloud Base 3660m, No Rain

Capella 22:27, 10 March 2017 (GMT) Checked latest image in R filter, peak value of 44,033 for target from file timestamped 22:18:25

Capella 22:30, 10 March 2017 (GMT) Checked latest image in G filter, peak value of 30,942 for target from file timestamped 22:22:08

Capella 22:30, 10 March 2017 (GMT) Checked latest image in B filter, peak value of 29,418 for target from file timestamped 22:23:09

Capella 22:59, 10 March 2017 (GMT) Sensor Check: Outside Temp 9.1; Outside Humidity 25%, Wind speed 16 kmph, Cloud Base 3630m, No Rain

Capella 22:59, 10 March 2017 (GMT) Checked latest image in R filter, peak value of 38,663 for target from file timestamped 22:52:18

Capella 22:59, 10 March 2017 (GMT) Checked latest image in G filter, peak value of 24,058 for target from file timestamped 22:44:03

Capella 22:59, 10 March 2017 (GMT) Checked latest image in B filter, peak value of 28,026 for target from file timestamped 22:48:30

Capella 23:07, 10 March 2017 (GMT) Contacted NDA to confirm team is happy to leave schedule to run for the evening.

11/03/2017

S382 PIRATE Project Deneb Group
Users: Janine Craig; Laura McCafferty; Brendan Quinn; John Forrest

Deneb 20:26, 11 March 2017 (GMT)Weather Conditions: Cloud Base too low (3150m). Dome closed. No observation possible at this time

Deneb 21:04, 11 March 2017 (GMT)Cloud base above required level for observation

Deneb 21:08, 11 March 2017 (GMT)Dome open now.  Preparing for test run
Test run R =20s, G=25s B=30s  Using Target J114749.04+351335.2 (KM Uma)

Deneb 21:22, 11 March 2017 (GMT)Test image showing extensive cloud cover

Deneb 21:24, 11 March 2017 (GMT)User:Deneb|Deneb]]Dome closed due to cloud base falling

Deneb 22:15, 11 March 2017 (GMT)Abandoned idea of using a test star due to adverse cloud cover conditions

Deneb 22:15, 11 March 2017 (GMT)Created Schedule Template ready for submission later. Aimed at target star. Exposures R= 20s; G=25s and B=35s. 5 repeats of each filter in sequence

Deneb 22:58, 11 March 2017 (GMT)Dome opening - cloud base rising. Preparing to submit test run on target object

Test run Exposures R= 20s; G=25s and B=35s.  Single run no repeats

Deneb 23:01, 11 March 2017 (GMT)Test image poor. Repeating test

Deneb 23:13, 11 March 2017 (GMT)Test results analysed. Low exposure due to cloud cover. For R filter counts=8364; G=10,516 counts ; B=8501 counts

Revised test to increase exposures to R= 30s; G=35s and B=50s.

Deneb 23:30, 11 March 2017 (GMT)Test submitted

Agreed final schedule with exposures of R= 20s; G=25s and B=35s.  5 repeats of each filter in sequence

Deneb 00:01, 12 March 2017 (GMT)Overnight observation schedule submitted

Deneb 00:07, 12 March 2017 (GMT)All logged off

13/03/2017

S382 PIRATE Project Vega Group
Users : Matt Williams, Sam Arnsby, Gavin Thrippleton 

Vega 19:35, 13 March 2017 (GMT) Target Star J13 28 37.5 +35 33 11.6

Vega 19:38, 13 March 2017 (GMT) logger : Matt

Vega 19:38, 13 March 2017 (GMT) observer : Sam

Vega 19:38, 13 March 2017 (GMT) monitor : Gavin

Vega 19:42, 13 March 2017 (GMT) Outside temp 4.86⁰C

Vega 19:42, 13 March 2017 (GMT) Outside humidity 30.07%

Vega 19:42, 13 March 2017 (GMT) Wind speed 19.30km/h, peaking at 27.54km/h, not raining

Vega 20:07, 13 March 2017 (GMT) Sample Test Red 52s, Green 55s, Blue 80s, 3 repeats on sample star RA 5 17 40.27 / DEC 33 53 55.39

Vega 20:26, 13 March 2017 (GMT) Sample Test Red peak 16,558, Green peak 15,797, Blue peak 21,247

Vega 20:46, 13 March 2017 (GMT) Sample Test Red 72s, Green 75s, Blue 100s, 3 repeats on sample star RA 5 17 40.27 / DEC 33 53 55.39

Vega 20:55, 13 March 2017 (GMT) Sample Test Red peak 26,003, Green peak 33,324, Blue peak 29,171

Vega 21:00, 13 March 2017 (GMT) Sample Test Red 60s, Green 63s, Blue 90s, 3 repeats on sample star RA 5 17 40.27 / DEC 33 53 55.39

Vega 21:12, 13 March 2017 (GMT) Sample Test Red peak 30,329, Green peak 27,141, Blue peak 29,788

Vega 21:15, 13 March 2017 (GMT) Sample Test Red 42s, Green 45s, Blue 70s, 3 repeats on sample star RA 5 17 40.27 / DEC 33 53 55.39

Vega 21:25, 13 March 2017 (GMT) Sample Test Red peak 20,937, Green peak 23,105, Blue peak 26,118

Vega 21:36, 13 March 2017 (GMT) Sample Test Red 15s, Green 20s, Blue 25s, 10s Luminance, 1 repeat on sample object de Mairan's Nebula (M43)

Vega 21:43, 13 March 2017 (GMT) Sample Test Red 6s, Green 12s, Blue 16s, 4s Luminance, 1 repeat on sample object de Mairan's Nebula (M43)

Vega 21:54, 13 March 2017 (GMT) Sample Test Red 2s, Green 4s, Blue 6s, 2s Luminance, 1 repeat on sample object Reflection Nebula (M78)

Vega 21:57, 13 March 2017 (GMT) Sample Test Red 5s, Green 10s, Blue 15s, 5s Luminance, 1 repeat on sample object Reflection Nebula (M78)

Vega 22:03, 13 March 2017 (GMT) Sample Test Red 2s, Green 4s, Blue 6s, 2s Luminance, 1 repeat on sample object Pleiades Cluster (M45)

Vega 22:09, 13 March 2017 (GMT) Sample Test Red 3s, Green 6s, Blue 9s, 3s Luminance, 1 repeat on sample object Beehive Cluster (M44)

Vega 22:12, 13 March 2017 (GMT) Sample Test Red 3s, Green 6s, Blue 9s, 3s Luminance, 1 repeat on sample object Beehive Cluster (M44) (repeat measure)

Vega 22:15, 13 March 2017 (GMT) Sample Test Red 2s, Green 4s, Blue 6s, 2s Luminance, 1 repeat on sample object Crab Nebula (M1)

Vega 22:26, 13 March 2017 (GMT) Sample Test Red 6s, Green 12s, Blue 16s, 4s Luminance, 1 repeat on sample object de Mairan's Nebula (M43) (repeat measure from 21:43)

Vega 22:26, 13 March 2017 (GMT) Sample Test Red 6s, Green 12s, Blue 16s, 4s Luminance, 10 repeats on sample object de Mairan's Nebula (M43)

Vega 22:33, 13 March 2017 (GMT) Sample Test Red 52s, Green 55s, Blue 80s, 1 repeat on Target Star J13 28 37.5 +35 33 11.6 (Test)

Vega 22:45, 13 March 2017 (GMT) Sample Test Red 45s, Green 50s, Blue 70s, 1 repeat on Target Star J13 28 37.5 +35 33 11.6 (Test)

Vega 22:51, 13 March 2017 (GMT) Sample Test Red 45s, Green 47s, Blue 60s, 1 repeat on Target Star J13 28 37.5 +35 33 11.6 (Test)

Vega 23:16, 13 March 2017 (GMT) Data collection Red 45s, Green 47s, Blue 60s, 3 objects observations repeat from 23:16 to 0624 on Target Star J13 28 37.5 +35 33 11.6

Vega 00:01, 13 March 2017 (GMT) Data collection Red 45s, Green 47s, Blue 60s, 3 objects observations repeat from 23:16 to 0624 on Target Star J13 28 37.5 +35 33 11.6 from before have been checked and scheduled run left overnight.

14/03/2017

S382 PIRATE Project Alpheratz Group
Users : Mark Woodland, Matthew Edgar, Miguel angel Andrades
Alpheratz 19:16, 14 March 2017 (GMT) Main Target Star : J114749.04+351335.2


Alpheratz 19:16, 14 March 2017 (GMT) Logger : Miguel

Alpheratz 19:16, 14 March 2017 (GMT) observer : Mark

Alpheratz 19:16, 14 March 2017 (GMT) monitor : Matt

Alpheratz 19:25, 14 March 2017 (GMT) deciding upon a target while waiting for our target to rise above 30degrees

Alpheratz 19:27, 14 March 2017 (GMT) M43 agreed upon

Alpheratz 19:37, 14 March 2017 (GMT) Outside temp 6.82⁰C, Outside humidity 0.98%, all other ambient factors are also within limits

Alpheratz 19:37, 14 March 2017 (GMT) Filter: Luminacy, exposure 30s, binning 2x2, repeat 1

Alpheratz 19:58, 14 March 2017 (GMT) Night time, observing starts.

Alpheratz 20:02, 14 March 2017 (GMT) First image clear but too bright, exposure reduced to 15s

Alpheratz 20:05, 14 March 2017 (GMT) Trying alpha III, 30s exposure

Alpheratz 20:09, 14 March 2017 (GMT) RGB 1 frame of each, 15s exposure

Alpheratz 20:15, 14 March 2017 (GMT) setting coordinates for M44 Filter: Luminacy, exposure 30s, binning 2x2, repeat 1

Alpheratz 20:31, 14 March 2017 (GMT) creating simple schedule for test R=9s, G=11s, B=18s

Alpheratz 20:31, 14 March 2017 (GMT) immediate queue acknowledge failure shown, NDA takes over to try to clear it.

Alpheratz 20:35, 14 March 2017 (GMT) fixed

Alpheratz 20:40, 14 March 2017 (GMT) command waiting transition failed on PIRATE, NDA to look into it again.

Alpheratz 21:17, 14 March 2017 (GMT) NDA fixed the problem, communication between the control software and telescope malfunctioned and the software had to be restarted. After this, the telescope control was returned to us. Retrying same parameters

Alpheratz 21:31, 14 March 2017 (GMT) target very faint, retrying same parameters but with R=11s, G=13s, B=20s.

Alpheratz 22:00, 14 March 2017 (GMT) main sequence submitted with R=11s, G=13s, B=20s, start time 20.25- end time 06.15, repeats of 1 for each filter 3 times with a 2x2 binning.

Alpheratz 22:00, 14 March 2017 (GMT) we are advised to stay around and check a few images before heading out.

15/03/2017

Capella 19:13, 15 March 2017 (GMT) S382 PIRATE PROJECT CAPELLA GROUP

Users: Howard Mitchell, Courtney Williams, Joseph McKay, Billy Dawborn, John Phelan.
Initial roles: Joe - observer's log, Billy - telescope, Courtney/John/Howard - environmental/webcam

Capella 19:18, 15 March 2017 (GMT) Howard questioned Fiona about our intentions for this evening.

Capella 19:20, 15 March 2017 (GMT) Fiona gave the team some ideas about taking pretty pictures of nebulae.

Capella 19:21, 15 March 2017 (GMT) Fiona quizzed the group about our preparations for tonight's session Capella 19:23, 15 March 2017 (GMT) 19:23, 15 March 2017 (GMT) A short discussion of nebulae followed with Howard telling the group of recommended magnitudes to observe

Capella 19:25, 15 March 2017 (GMT) An outbreak of morale.

Capella 19:26, 15 March 2017 (GMT)The group discussed the forthcoming weather and tried to foresee any difficulties.

Capella 19:30, 15 March 2017 (GMT)Weather report. Outside temperature 2.57 degrees C, Humidity 44.11%, Wind Speed 32 kmh, Cloud Base 3685 m, Rain - dry, Dome - open, temperature -20 Degrees C, cooled.

Capella 19:31, 15 March 2017 (GMT) Team adding schedules.

Capella 19:34, 15 March 2017 (GMT) Target - Crab Nebula - and the group discussing a schedule for observing it.

Capella 19:36, 15 March 2017 (GMT) Billy discussing his schedule with Howard.

Capella 19:38, 15 March 2017 (GMT) Luminance, billing and exposures discussed and decided upon.

Capella 19:41, 15 March 2017 (GMT) Howard suggested changing filters to Red, Green and Blue for the next exposures.

Capella 19:45, 15 March 2017 (GMT) Billy having trouble with the program being difficult.

Capella 19:46, 15 March 2017 (GMT) NDA checking the system for us.

Capella 19:58, 15 March 2017 (GMT) New schedule given by NDA

Capella 20:05, 15 March 2017 (GMT) Schedule for Crab Nebula submitted

Capella 20:06, 15 March 2017 (GMT) Billy still having problems with the schedule

Capella 20:09, 15 March 2017 (GMT) NDA asked for help with the scheduling program.

Capella 20:13, 15 March 2017 (GMT) Billy completing the schedule box, but the program is not working well

Capella 20:18, 15 March 2017 (GMT) Crab nebula finally imaged

Capella 20:19, 15 March 2017 (GMT) being downloaded by Howard for a count.

Capella 20:21, 15 March 2017 (GMT) Courtney now taking on the scheduling role alongside Billy.

Capella 20:23, 15 March 2017 (GMT) Crab Nebula 2 being imaged

Capella 20:28, 15 March 2017 (GMT) Roles switched (Courtney telescope, Billy observers log, John/Howard/Joseph Environmental monitor/webcam).

Capella 20:30, 15 March 2017 (GMT) Scheduled submitted for Crab nebula. 3 images, no filter (luminance) 60 second exposure 2x2 binning.

Capella 20:33, 15 March 2017 (GMT) Images checked for count rate. Reading at about 2000 counts (60 second exposure).

Capella 20:38, 15 March 2017 (GMT) Submitted schedule for crab nebula. 3 minute exposure in luminance, red, green and blue filters. 1 image in each filter.

Capella 20:51, 15 March 2017 (GMT) Checked luminance image in AstroimageJ. Image has better detail but the middle is more blurred than 60 second exposure.

Capella 20:57, 15 March 2017 (GMT) Switched roles: Courtney - log, Joe - telescope, John/Billy/Howard - environmental monitor/webcam

Capella 20:58, 15 March 2017 (GMT) Joe is submitting the test schedule for our target star (R - 50s, G - 50s, B - 100s; one repeat of each, 2x2 binning)

Capella 21:04, 15 March 2017 (GMT) Joe is clearing the queue and submitting

Capella 21:08, 15 March 2017 (GMT) Needed to re-clear the queue and re-submit, now the test run is going

Capella 21:13, 15 March 2017 (GMT) Test frames look good, count rate as expected, so John is now taking over as telescope operator to set up the schedule

Capella 21:18, 15 March 2017 (GMT) John is setting up the schedule for this night's main observing run

Capella 21:19, 15 March 2017 (GMT) Ensuring we set the repeat as 1 to avoid having lots of files to rename

Capella 21:21, 15 March 2017 (GMT) Other team members checking the schedule

Capella 21:24, 15 March 2017 (GMT) Deleted the G and B runs from our schedule as we aren't doing these

Capella 21:28, 15 March 2017 (GMT) Everyone has checked the schedule and is happy with it - though we are changing start time to 2140 rather than 2200

Capella 21:30, 15 March 2017 (GMT) Environmental check: 0.67 deg C, outside humidity 50.71%, wind speed 18.98 km/h, cloud base 3504.56m, not raining, dome open, camera at -20 deg C (all sensors within limits)

Capella 21:31, 15 March 2017 (GMT) John has cleared the queue and submitted the schedule

Capella 21:35, 15 March 2017 (GMT) We will wait for the first image to show up in Astrodrive to check all is well

Capella 21:43, 15 March 2017 (GMT) We are looking at the first image in AstroImageJ

Capella 21:44, 15 March 2017 (GMT) Joe left for the evening

Capella 21:51, 15 March 2017 (GMT) Howard will put together the zip files of our images tomorrow morning

Capella 22:02, 15 March 2017 (GMT) Appears that we have just passed the peak - count rate increased, now going back down

Capella 22:07, 15 March 2017 (GMT) We are checking the counts to ensure we are all getting the same results

Capella 22:08, 15 March 2017 (GMT) Dome closed due to cloud base and humidity

Capella 22:12, 15 March 2017 (GMT) NDA advised us to leave it as we can't do anything else - hopefully the weather will clear up

Capella 22:15, 15 March 2017 (GMT) Cloud base is dropping fast!

Capella 22:21, 15 March 2017 (GMT) Everyone has exited Abot WITHOUT LOGGING OUT! Observing ended for the night - the cloud was rising again so hopefully we will get some results. NDA has said he will check the programme around midnight just in case.

Capella 22:25, 15 March 2017 (GMT) Skype call ended - Howard doing pixel shift test for the NDA.

16/03/2017

S382 PIRATE Project Deneb Group

Users: Daniel Shaw; Terry Welsh; Sean Hunt; Janine Craig
Roles: Daniel - Lead; Terry - Analyst; Sean - Log; Janine - Environment

Deneb 20:04, 16 March 2017 (GMT) Bad weather meaning the prospect of getting any data not looking promising

Deneb 20:04, 16 March 2017 (GMT) Assessing whether previous exposure times should be adjusted

Deneb 20:04, 16 March 2017 (GMT) Checked frame with blue filter at peak, which fell in suitable range, kept previous exposure time of 35s

Deneb 20:04, 16 March 2017 (GMT) Checked frame with red filter at peak, which fell in suitable range, kept previous exposure time of 20s

Deneb 20:04, 16 March 2017 (GMT) Checked frame with green filter at peak, which fell in suitable range, kept previous exposure time of 25s

Deneb 20:14, 16 March 2017 (GMT) Set up exposure schedule in the hopes of improvement in the weather

Deneb 20:31, 16 March 2017 (GMT) Still raining

Deneb 20:40, 16 March 2017 (GMT) After checking further frames, decided to add 5 seconds to all exposure times

Deneb 20:40, 16 March 2017 (GMT) It's also still raining

Deneb 20:46, 16 March 2017 (GMT) Submitted schedule with each filter to repeat 5 times

Deneb 20:46, 16 March 2017 (GMT) The rain continues

Deneb 21:39, 16 March 2017 (GMT) Weather conditions have gotten worse and NDA has suggested we call it a night

Deneb 21:39, 16 March 2017 (GMT) All logged off

17/03/2017

S382 PIRATE Project Pollux Group

Target: 1SWASP J132837.25+353311.6

Users: Simon Bowyer, Billy Evans, Simon Bell, Sami Zenasni, Victoria Stanbridge

Pollux 20:24, 17 March 2017 (GMT) Weather not permitting observation, time delay used to discuss observation schedule

Pollux 20:30, 17 March 2017 (GMT) Observation schedule submitted

Pollux 20:40, 17 March 2017 (GMT) Immediate queue cleared

Pollux 20:50, 17 March 2017 (GMT) Still raining

Pollux 21:53, 17 March 2017 (GMT) Still raining

Pollux 22:35, 17 March 2017 (GMT) Still raining, all logged off as unlikely rain will stop, if it does observing schedule has been submitted.

18/03/2017

S382 PIRATE Project Arcturus Group

Target: 1SWASP J112541.63+423449.6 (BS UMa)

Users: Sean Moran, Richard Ballard, Alex Davidson, Daniel Toppin, Vincent Tully.

Roles: Log Keeper - Alex D, PIRATE Operator - Sean M/Richard B

Arcturus 19:37, 18 March 2017 (GMT) Team joined skype call. Initial weather readings show heavy rain and fog

Arcturus 19:48, 18 March 2017 (GMT) Briefing with Fraser. Discussing weather and unlikelihood of success this evening. Target visible after 8pm. Discussion about whether we should use a clear filter for taking images. Decided it is not necessary due to the capabilities of PIRATE.

Arcturus 19:52, 18 March 2017 (GMT) Discussed previous exposure times and filters used (R = 60s, G = 60s, B = 120s)

Arcturus 20:01, 18 March 2017 (GMT) Discussed likelihood of rain continuing all night, decided that a schedule would be set up to run with the assumption that the rain would clear - just in case. Fraser logged off and left us to discuss our observing strategy.

Arcturus 20:14, 18 March 2017 (GMT) Discussed filter repeats options, whether we will run each filter 3 times before moving on or if we will take 1 image with each filter before changing them. The team seemed happy to go with the initial set up of 3x repeat (RRR-GGG-BBB).

Arcturus 20:16, 18 March 2017 (GMT) Still raining. Outside temp just below freezing and humidity around 99%.

Arcturus 20:33, 18 March 2017 (GMT) Due to incomplete knowledge of the filter wheel and the order in which the filters and loaded and changed, the team decided to go with R>G>B order since this has been proven to be successful during the first session on 9th March. If more information was available we could adjust this order but the team is happy as is.

Arcturus 20:33, 18 March 2017 (GMT) Richard is currently working through the instructions to create the long run schedule.

Arcturus 20:40, 18 March 2017 (GMT) Richard has created the template and is filling in the details along with Sean who has opened the previous night's template to monitor the data entry compared to the previous long-run schedule.

Arcturus 20:55, 18 March 2017 (GMT) Long-run schedule added and upon review by the team it was noticed that the B filter didn't add to the schedule. Richard edited the program and realised that the "add" button wasn't clicked. Team reviewed the schedule afterwards are were happy with the final form.

Arcturus 21:03, 18 March 2017 (GMT) NDA gave green light to submit schedule and he will review it. Slight issue with getting the immediate queue cleared due to refresh speed of queue, it repopulated quickly after initiating the clear command.

Arcturus 21:10, 18 March 2017 (GMT) NDA had to take control of PIRATE to sort queue issues. Immediate queue now cleared successfully and the long-run schedule has been successfully submitted. NDA confirmed that schedule looked good and everything was okay. PIRATE is now tracking.

Arcturus 21:10, 18 March 2017 (GMT) Still raining. NDA raised a slight concern that PIRATE shouldn't track while the dome is closed. PIRATE Operator turned on IR light in order for NDA to check camera movement. NDA has taken control to stop tracking as this shouldn't be happening.

Arcturus 21:17, 18 March 2017 (GMT) Team noted that rain stopped around midnight for previous night so we are hopefull this happens again. At the moment the rain doesn't look to be clearing though.

Arcturus 21:23, 18 March 2017 (GMT) Issues with taking control of PIRATE. Telescope accepted the re-submitted schedule but our Operator did not have control. NDA looking into it. Another team member took control of PIRATE (S Moran) due to possible webpage/internet issues with Richard.

Arcturus 21:23, 18 March 2017 (GMT) NDA has confirmed that PIRATE is now set to take our images if the rain clears.

Arcturus 21:25, 18 March 2017 (GMT) PIRATE Operator has closed the browser and we are happy with the schedule. Rain is still heavy.

Arcturus 21:31, 18 March 2017 (GMT) Team chatting about data / lightcurves from previous data and the photometry process. Related literature for the target star discussed.

Arcturus 21:39, 18 March 2017 (GMT) Team ended skype call and logged off for the night.

20/03/2017

Alperatz PIRATE group S382

Observers: Darren Smith, Gavin Duck, Jonathan Ritchie, Lester Sandles, Matthew Edgar
PIRATE operator: Gavin. Log keeper: Jonathan. Environment/data monitors: Darren, Lester, Matt.

Alpheratz 20:26, 20 March 2017 (GMT) Targetting M81 whilst waiting for KM UMa to be visible.

Alpheratz 20:26, 20 March 2017 (GMT) Outside humidity 23.3%

Alpheratz 20:26, 20 March 2017 (GMT) Connection issues with the PIRATE dashboard, refreshing browsers seems to correct this.

Alpheratz 20:31, 20 March 2017 (GMT) Weather conditions all still within limits.

Alpheratz 20:47, 20 March 2017 (GMT) Luminance filter is giving peak count of 3K at 2 sec exposure.

Alpheratz 20:47, 20 March 2017 (GMT) Increasing exposure times for luminance, red, green and blue filters.

Alpheratz 20:47, 20 March 2017 (GMT) Luminance 10 sec, red 15 sec, green 20 sec, blue 25 sec.

Alpheratz 20:47, 20 March 2017 (GMT) Good image quality for M81.

Alpheratz 20:47, 20 March 2017 (GMT) Luminance 30 sec, red 60 sec, green 100, blue 125 sec.

Alpheratz 21:08, 20 March 2017 (GMT) PIRATE connection problems for some group members again. Dashboard says 'Open University is disconnected'.

Alpheratz 21:08, 20 March 2017 (GMT) Image quality being assessed with AstroImageJ.

Alpheratz 21:08, 20 March 2017 (GMT) Outside humidity 12.98%.

Alpheratz 21:08, 20 March 2017 (GMT) Final attempt at M81. Luminance 45 sec, red 80 sec, green 100 sec, blue 200 sec.

Moving on to target star: KM UMa, J114749.04+351335.2

Alpheratz 21:08, 20 March 2017 (GMT) Exposure times: red 11 sec, green 13 sec, blue 20 sec.

Alpheratz 21:13, 20 March 2017 (GMT) Test images in progress.

Alpheratz 21:13, 20 March 2017 (GMT) Weather conditions all okay.

Alpheratz 21:33, 20 March 2017 (GMT) Outside humidity 6.40%

Alpheratz 21:33, 20 March 2017 (GMT) Test images being checked on Astrodrive. Peak counts all looking acceptable for R, G and B filters.

Alpheratz 21:33, 20 March 2017 (GMT) Schedule being set up for automated session.

Alpheratz 21:39, 20 March 2017 (GMT) Schedule needs editing.

Alpheratz 21:39, 20 March 2017 (GMT) Schedule submitted.

Alpheratz 21:44, 20 March 2017 (GMT) Date correction on schedule. Resubmitted.

Alpheratz 21:44, 20 March 2017 (GMT) Schedule checked and all okay.

Alpheratz 21:44, 20 March 2017 (GMT) Initial images being checked.

Alpheratz 22:07, 20 March 2017 (GMT) Group logging off PIRATE.

21/03/2017

S382 PIRATE Project Vega Group

Target: J132837.25+353311.6

Users: Theron Pearman, Fleur Perry, John Marshall

Roles: Log Keeper - John, PIRATE Operator - Theron / Fleur

Vega 20:44, 21 March 2017 (GMT) Images of M37 whilst awaiting Target

Vega 20:46, 21 March 2017 (GMT) Outside temperature 5.31 C

Vega 20:46, 21 March 2017 (GMT) Outside Humidity 7.96%

Vega 20:46, 21 March 2017 (GMT) All sky camera - mostly green

Vega 20:51, 21 March 2017 (GMT) Images of M42

Vega 20:51, 21 March 2017 (GMT) Intermittent connection issues

Vega 21:20, 21 March 2017 (GMT) Images of M97 - The Owl Nebula

Vega 22:00, 21 March 2017 (GMT) Target test 1 60s Blue, 47s Green, 45s Red

Vega 22:20, 21 March 2017 (GMT) Target test 2 50s Blue, 37s Green, 35s Red

Vega 22:29, 21 March 2017 (GMT) Target test 3 45s Blue, 32s Green, 30s Red

Vega 22:46, 21 March 2017 (GMT) Outside temperature 6.71 C

Vega 22:46, 21 March 2017 (GMT) Outside Humidity 10.79%

Vega 22:46, 21 March 2017 (GMT) All sky camera - mostly green

Vega 22:47, 21 March 2017 (GMT) Target live run 45s Blue, 32s Green, 30s Red

Vega 22:49, 21 March 2017 (GMT) Target live run - Laid in and started

Vega 22:59, 21 March 2017 (GMT) Initial images look very nice

Vega 22:59, 21 March 2017 (GMT) Automatic program set in

Vega 23:11, 21 March 2017 (GMT) And so to bed

22/03/2017

Arcturus PIRATE group S382

Observers: Courtney Goddard, Daniel Sharpley, Andy Bettger, Sean Moran, Alex Roberts

PIRATE operator: Andy Log keeper: Courtney

Target: 1SWASP J112541.63+423449.6 11h25m41.0s 42°34'49.0"

Arcturus 20:15 Weather favourable, possible cloud.

Arcturus 20:15 Testing at R 60s G 60s B 120s

Arcturus 20:20 initial counts returned at R 7469 G 5459 B 4967

Arcturus 20:23 increasing exposure time 50% to Red 90s Green 90s Blue 180s

Arcturus 20:30 counts returned 4353 5904 5764 (possible cloud cover)

Arcturus 20:34 red counts returned 7384 7465 7690

Arcturus 20:40 green counts returned 6333 6171

Arcturus 20:41 consensus to leave exposure at R 90s G 90s B 180s

Arcturus 20:43 blue counts increased to 7309 from 180s exposure.

Arcturus 20:57 Long schedule set "Long_Run_Schedule_Arcturus_20170322" 3x R 90s 3x G 90s 3x B 180s

Arcturus 21:10 counts upto 18000.

23/03/2017

Pollux 20:01, 23 March 2017 (GMT) Logger: Stevie
Pollux 20:06, 23 March 2017 (GMT) Environment: Andy
Pollux 20:06, 23 March 2017 (GMT) Operator: Adam
Pollux 20:16, 23 March 2017 (GMT) Webcams: Billy
Pollux 20:16, 23 March 2017 (GMT) Data analysis: Victoria and Simon
Pollux 20:17, 23 March 2017 (GMT) M42 viewed as a test, image came out too bright 20 sec exposure RGB filter
Pollux 20:27, 23 March 2017 (GMT) Trying again hydrogen alpha filter 30 sec exposure
Pollux 20:27, 23 March 2017 (GMT) moving on to horse-head nebula
Pollux 20:35, 23 March 2017 (GMT) Webcams all look fine
Pollux 20:35, 23 March 2017 (GMT) telescope has moved to horse-head nebula
Pollux 20:43, 23 March 2017 (GMT) Image too dark doubling exposure time to 40 sec
Pollux 20:43, 23 March 2017 (GMT) Better images in the red filter exposure upped to 80 sec
Pollux 20:50, 23 March 2017 (GMT) Moving on to M81 exposure time 20s
Pollux 20:50, 23 March 2017 (GMT) M81 exposure increased to 80s
Pollux 20:57, 23 March 2017 (GMT) Red filter M81 100 sec exposure
Pollux 21:03, 23 March 2017 (GMT) Moving to M44 RGB 20s exposure time
Pollux 21:14, 23 March 2017 (GMT) Moved to M82 with RGB filter 20s exposure time
Pollux 21:28, 23 March 2017 (GMT) Image shows signs of M82 exposure time increased to 80s in hopes of getting a better image
Pollux 21:28, 23 March 2017 (GMT) Template for target star J132837.25+353311.6 has been input for test exposure in RGB broadband filters with the following exposure times R= 90s G= 100s B= 120s
Pollux 21:48, 23 March 2017 (GMT) Test exposure started
Pollux 21:48, 23 March 2017 (GMT) 15000 counts in test red image
Pollux 21:48, 23 March 2017 (GMT) 11700 counts in test green image
Pollux 21:48, 23 March 2017 (GMT) 12000 counts in test blue
Pollux 21:59, 23 March 2017 (GMT) Long run started R= 90s G and B= 120s

Weather parameters

Outside temperature
Within limits

2.73°C
21:00 UTC 21:30 UTC 22:00 UTC
1.87 °C    2.91 °C  3.94 °C 4.97 °C 6.01 °C 6.01 °C
Outside Humidity
Within limits

19.72%
Wind speed
Within limits

23.82km/h
Cloud base
Within limits
4081.18m
Pollux 22:11, 23 March 2017 (GMT) Repeats changed from Rx3 to Rx1 Rx1 Rx1

28/03/2017

S382 PIRATE Project Deneb Group
Observers: Daniel Shaw; Terry Welsh; Brendan Quinn

Deneb 19:40, 28 March 2017 (UTC)Weather Conditions: Fine and within limits. Dome Open. Pirate controller: Daniel Shaw Analysis: Terry Welsh log: Daniel Shaw Weather: Terry Welsh

Deneb 19:45, 28 March 2017 (UTC) Preparation completed; advised by PNDA not to submit anything until night time (20:06 UTC). Current parameters for J114749.04+351335.2 are R25, G30, B40 on a 2x2 binning.

Deneb 19:53, 28 March 2017 (UTC) Dome closed due to calibration completion.

Deneb 20:07, 28 March 2017 (UTC) Dome open, schedule submitted.

Deneb 20:20, 28 March 2017 (UTC) Images not showing up in AstroDrive; PNDA recommends clearing queue and resubmitting.

Deneb 20:21, 28 March 2017 (UTC) Schedule cleared and resubmitted. Current parameters for J114749.04+351335.2 are R25, G30, B40 on a 2x2 binning.

Deneb 20:25, 28 March 2017 (UTC) Brendan Quinn joined group, takes over as log writer.

Deneb 20:30, 28 March 2017 (UTC) NDA advised that images were being saved on AstroDrive but were not currently visible.

Deneb 20:44, 28 March 2017 (UTC) Test results: Red peak 55432, Green 39880, Blue 53346 (all too high).

Deneb 20:48, 28 March 2017 (UTC) Adjusting exposure times to rerun tests with R20, G30, B32.

Deneb 20:53, 28 March 2017 (UTC) Adjourned for short time to allow software developers to fix AstroDrive problem.

Deneb 21:05, 28 March 2017 (UTC) Resumed session after AstroDrive fix and seems to be working now.

Deneb 21:06, 28 March 2017 (UTC) Submitting tests with R20, G30, B32.

Deneb 21:20, 28 March 2017 (UTC) Results Red 54269, Green 65535, Blue 65535. All much too high and not as expected - 65535 is max value possible. Star expected to be at 50 degrees now and at zenith approx 00:30.

Deneb 21:31, 28 March 2017 (UTC) Submitting tests with R10, G15, B16.

Deneb 21:41, 28 March 2017 (UTC) Results Red 22065, Green 33012, Blue 27230. G, B too high, aiming for around 20000 each to avoid overexposure at zenith of target star.

Deneb 21:52, 28 March 2017 (UTC) Submitting tests with R10, G10, B12.

Deneb 22:10, 28 March 2017 (UTC) Results Red 19971, Green 17371, Blue 16577. These are in range 10k-20k which seems OK.

Deneb 22:19, 28 March 2017 (UTC) Submitting overnight run with most recent test exposures R10, G10, B12, 5 repeats for each filter.

Deneb 22:20, 28 March 2017 (UTC) Conditions looking good for overnight run. Cloudbase 3944m. Humidity 16.27%.

Deneb 22:30, 28 March 2017 (UTC) Results of first batch Red 15882, Green 16346, Blue 13717.

Deneb 22:41, 28 March 2017 (UTC) Group logging off.

29/03/2017=

S382 PIRATE Project Pollux Group

Target: 1SWASP J132837.25+353311.6

Users: Simon Bell, Sami Zenasni, Victoria Stanbridge

Pollux 21:03, 29 March 2017 (BST) All users logged onto ABOT

Pollux 21:03, 29 March 2017 (BST) All weather conditions within acceptable limits

Pollux 21:03, 29 March 2017 (BST) Waiting for the dome to open

Pollux 23:00, 29 March 2017 (BST) Target not visible until approximately 22:15, using time to image various regions of the sky

Pollux 23:00, 29 March 2017 (BST) First images of target taken, currently not above 30 degrees, checking photon count

Pollux 23:00, 29 March 2017 (BST) Reducing exposure times on R and G as photon counts too high; B 120; G 100; R 70

Pollux 23:00, 29 March 2017 (BST) Reducing exposure time again as photon counts still too high; B 100; G 80; R 50

Pollux 23:00, 29 March 2017 (BST) Reducing exposure time again as photon count still to high; G 70; R 40

Pollux 23:00, 29 March 2017 (BST) Reducing exposure time again as photon count still too high; G 50

Pollux 23:00, 29 March 2017 (BST) R filter is now okay at 40

Pollux 23:00, 29 March 2017 (BST) Reducing exposure time again as photon count still too high; G 40; B 80

Pollux 23:00, 29 March 2017 (BST) G filter is now okay at 40

Pollux 23:49, 29 March 2017 (BST) There are issues with the naming of files, files on AstroDrive taken in the B filter have 'G' in the title not 'B'

Pollux 23:49, 29 March 2017 (BST) Reducing exposure time on B to 60

Pollux 23:49, 29 March 2017 (BST) Exposure times and photon counts now okay, NDA investigating 'naming' issues

Pollux 23:49, 29 March 2017 (BST) Observing schedule submitted

Pollux 23:53, 29 March 2017 (BST) Observing schedule name updated and re-submitted to avoid confusion with first 69 images taken

Pollux 23:53, 29 March 2017 (BST) All users logged off


30/03/2017=

S382 PIRATE Project Arcturus Group

Target: BS Uma

Users: Sean Moran, Alex Davidson, Richard Ballard

Arcturus 20:23, 30 March 2017 (BST) Team logged onto Skype. Awaiting Richard. Weather is good, slight cloud. Dome opened at 8:20pm. Richard takes over log Alex setting up exposure for long schedule in Red Arcturus 20:48, 30 March 2017 (BST)Discussion for setting up 'Red 1 repeat' schedule

Arcturus 20:53, 30 March 2017 (BST)Alex D has control of Pirate

Arcturus 20:54, 30 March 2017 (BST)Sean M notes dome has just closed

Arcturus 20:57, 30 March 2017 (BST)outside temperature 10.70 degrees, not raining

Arcturus 21:00, 30 March 2017 (BST)Sean M and Alex D working on Long_Run_Arcturus_2017_03_30 schedule

Arcturus 21:05, 30 March 2017 (BST)Alex D notes that he can now take control of Pirate

Arcturus 21:06, 30 March 2017 (BST) Alex D notes that dome may be about to open

Arcturus 21:09, 30 March 2017 (BST) NDA agrees that sighting shots can now be taken

Arcturus 21:09, 30 March 2017 (BST) Dome open

Arcturus 21:11, 30 March 2017 (BST) Image acquired

Arcturus 21:12, 30 March 2017 (BST) Image appears in Astrodrive and Alex D is downloading it to check quality

Arcturus 21:15, 30 March 2017 (BST) Alex D notes peak count of 18,000

Arcturus 21:15, 30 March 2017 (BST) NDA steps away awaiting submission of schedule and/or any problems

Arcturus 21:15, 30 March 2017 (BST) cloud moving in from West

Arcturus 21:16, 30 March 2017 (BST)Sean M checks the long schedule, and 'end time' has been selected for longest available 'run'

Arcturus 21:20, 30 March 2017 (BST) Sean notes 'all sky' showing clear at zenith

Arcturus 21:20, 30 March 2017 (BST) Sean/Alex/Richard have checked the schedule: 60s in R only

Arcturus 21:25, 30 March 2017 (BST)NDA is checking target and time (UTC) at which it will achieve zenith

Arcturus 21:25, 30 March 2017 (BST) NDA checks RA and DEC of target

Arcturus 21:25, 30 March 2017 (BST)NDA recommends 2000hrs commencement, and amendment for 0600 (rather than 0605) cessation

Arcturus 21:25, 30 March 2017 (BST)schedule submitted

Arcturus 21:29, 30 March 2017 (BST) Alex D changes start time of schedule to 2000 (UTC)..end time stays at 0600

Arcturus 21:30, 30 March 2017 (BST)NDA checks resubmitted schedule

Arcturus 21:42, 30 March 2017 (BST)NDA will be available for any problems, long run looks good

Arcturus 21:42, 30 March 2017 (BST)schedule in place

Arcturus 21:42, 30 March 2017 (BST) discussion about approach to data handling and photometery

Arcturus 21:45, 30 March 2017 (BST)initial frames being acquired, look good, and skype call ended


22/08/2017

SXPA288 User 22:14, 22 August 2017 (BST) 2017-08-22 Alan Cayless

SXPA288 User 22:18, 22 August 2017 (BST) Waiting for clouds to lift

SXPA288 User 22:34, 22 August 2017 (BST) Jupiter, Spica, Saturn visible but cloudbase still too low.

SXPA288 User 22:52, 22 August 2017 (BST) Clouds didn't lift.  All users logged off.

SXPA288 User 23:20, 22 August 2017 (BST) Logged back in. Clouds lifting. Queued acquisitions for M92 (B, G) and M27 (R, G, B)

23/08/2017

SXPA288 User 23:10, 23 August 2017 (BST) Queued Longer exposures of M27 (150 and 300 s)

SXPA288 User 23:17, 23 August 2017 (BST) Not all were successful - only R and G acquired.

SXPA288 User 23:17, 23 August 2017 (BST) B and one G listed in Completed Observations as Failed.


SXPA288 User 14:47, 13 September 2017 (BST) Test

22/09/2017

SXPA288 User 21:37, 22 September 2017 (BST) Pier Flip Test

SXPA288 User 21:37, 22 September 2017 (BST) M15 [E]

SXPA288 User 21:37, 22 September 2017 (BST) M13 [W]

SXPA288 User 21:37, 22 September 2017 (BST) M 2 [E]

SXPA288 User 21:37, 22 September 2017 (BST) M92 [W]

SXPA288 User 21:38, 22 September 2017 (BST) M56 [W]

SXPA288 User 21:38, 22 September 2017 (BST) M29 [E]

SXPA288 User 21:38, 22 September 2017 (BST) M39 [E]

SXPA288 User 21:43, 22 September 2017 (BST) M52 [E]